A Complete Bibliography of Publications in the C/C++ Users Journal: 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

12 August 2013 Version 1.07

Title word cross-reference

#1 [Hic00]. #2 [Cac00a]. #3 [Fos00]. #4 [Pet00]. #5 [Tim01]. #6 [Cha01]. #7 [Now01]. #8 [Abr02].

3 [Tel00]. N [Bav00b, Smi00]. X [Thi00]. Y [Thi00].

* [Now01].

- [Thi00]. -D [Tel00]. -Dimensional [Bav00b]. -Way [Smi00].

.NET [Lip02, Sch00d, Smi02].

0x [Sut02f].

20th [Ano01b]. 21st [Sch01a].

4 [Mey02c].

Accelerated [Gla01]. Access [Dix01, Kle00, Mor00a, Yor01]. Accident [KM01a]. across [Lop01]. Adaptable [Kem02b]. Adapter [May01]. Adapters [Dew02c, Dew02d]. Adaptive [Phi00c]. Adaptors [Bec02e, PW00]. Adding [Sta01]. Adventures [Sch00a]. Advice [Mey02a]. Alan [Sch01c]. Algol [Pla00a]. Algorithm [HR02, Lor01, Phi00a]. Algorithms [All00c, Ano02a, Bec02d, Buc02, Dew01c, Koe01, Mey01h, Mey01i]. Allocators [Vla02]. Alone [All02b]. Angelology [KB00a]. Annealing [Buc01]. Anniversary [Ano01b]. Application [Bro00, Tru02]. Applications [Mor00b, Mot02, Pad01, Ten02]. Applied [KB00c]. Applying [Abl02]. Approach [Käm01]. Apps [Ano02e]. Aquaman [Sch00a]. Arbitrary [Rin01]. Array [KM02a, Mey02c]. Arrays [All00a, Bav00b, Mey01e, Mey01f, Mey02d, Sak00a, Sak00b]. Art [All02a]. Ask [Sch01h]. Association [Til01]. Asynchronous [Nak01]. auto ptr [Vla01]. Automated [All00g]. Avoiding [KM02a]. Axis [Bra00].

Back [Bri01f, Dew01f]. Backtracking [Lab00]. Bad [Bec00h]. Ballad [Sch00k]. Based [Aus01c, Bra00, Cac00b, Fau00, Gui00c, Mor00a, Sca00, Smi00]. Basic [All00b]. Basics [Bec00e]. Be [Gri01]. Began [Mey00c]. Behavior [Yod01]. Behind [Vla01]. Best [All02g]. Better [Bec00i]. between [Lin01]. BGL [Abl02]. Big [Bri01e, Sut01b]. Binary [Hub02, KM01c]. Bit [Aus01a, Dew02a, Dew02b]. Bit-Wise [Dew02a, Dew02b]. Bits [Bec00f]. Bitsets [Aus01a]. Block [KB01a]. Blocks [Nas00, Que00]. Book [Sut02a, Bri00a, But00, Cir01, For01, Gla01, Käm01, Phi00b, Wal01]. Books [Ano01g]. bool [Sch02a, Mey02a]. Boost [Kar02]. Boost.Threads [Kem02a]. Both [All02g]. Boxes [Kem02b]. Boyer [Ber00]. Bridge [May01]. Brother [All02a, Bri01e]. Browser [Mor00a]. Browser-Based [Mor00a]. Buffers [Kue05]. Building [Bav00a, Hub01a].

C [Sut02a, Bri00a, But00, Cir01, For01, GF01, Gla01, Käm01, Pla00d, Pla00b, Pla00f, Pla00e, Pla00c, Sut02e, Sut02f, Zol01, Abr02, Ano01b, Ano01a, Ano01e, Ano02a, Ano02e, Bat00, Bav00a, Bec02a, Bec02b, BMG02, Bri00n, Cac00a, Cav02, Cha01, Col02, Cre00, Fau00, Fos00, Gat02, Gon00, Hic00, Koe01, KM01a, KM01d, KM01b, KM01e, KM01c, KM02d, KM02e, KM02a, KM02b, KM02c, KL02, LK00, Laf00, Lin01, Lip02, Liu01, Mey00b, Mey00c, Mey00a, Mey01a, Mey01b, Mey01c, Mey01d, Mey01e, Mey01f, Mey01g, Mey02a, Mey02b, Mey02c, Mor02, Nel01, Now01, Pet00, Pin02, Pla00e, Pla02, Rin01, Sak00a, Sak00b, Sew02, Smi02, Str02b, Str02a, Str02c, Sut01a, Tim01]. C# [Ano01g, Lip01]. C/C [Cir01, Pla00d, Pla00b, Pla00f, Pla00e, Pla00c, Abr02, Ano01b, Cac00a, Cha01, Fos00, Hic00, Lin01, Now01, Pet00, Tim01]. C/C [KL02, Pla02]. C99 [Mey00b, Mey00a]. Cafe [Sch00g]. Calculations [Bec00d]. Call [Hub01b]. Callbacks [Bat01]. Calling [Sew02]. Case [Str02b, Str02a]. Casts [Bec00a]. Catching [Gui00a]. CE [Gar00]. Century [Sch01a]. CGI [Mor00b]. Character [Sak00a, Sak00b]. Cheese [All02f]. choice [Sch02a]. Circles [Dew02e]. Class [Bav00b, Cha02, Ehl00, Fau00, Fox01, KM02b, Nik00, Pad01, Sew02, Swa00, Tru02]. Classes [All00e, Bec01d, Cal00, Cha01, Nie00, Pan01, Pee01, Que00, Rei00, Sme01, Til01]. Clipboard [Rog00a]. Code [Bon01, Geo02, Hub02]. Coder [KB01a]. Codes [Gui00a]. Collections [All00c, Kle00]. COM [Kou00, LK00, Mor00a]. Combinatorial [Lab00]. Combining [Ber00]. Comes [KB00e]. Command [Rog00a]. Comment [Tim01]. Common [Bec00b, Bec00c, Dew00, Dew01c, Dew01d, Dew01e, Dew01b, Dew01a, Dew01f, Dew02c, Dew02d, Dew02e]. comp.cuj.qa.moderated [Sch00b]. Compare [KB00d]. Comparisons [Bec00f]. Compatibility [Str02b, Str02a]. Compatible [Ano01f]. Compilation [Bec00l]. Compiler [Col02]. compilers [Sch02a]. Completely [Aus02a]. Complex [Aus02b]. Complexity [Bri01j]. Component [Gui00c]. Component-Based [Gui00c]. Components [Bia00, May01]. Compound [Mey01a]. Compressed [Yor01]. Computational [Abl02]. Concepts [Vla01]. Configuration [Bro00]. Conformance [Pla00f, Pla00e, Sut01a]. Const [Aus01b]. Constants [Cac00a]. Constructed [Mor02]. Constructor [Sut00]. Container [Bec01b, KM01e, PW00]. Containers [Aus01c, Bec01c, Sut01b]. content [Sch02a]. Contour [Wu00]. Contributing [Cir01]. Control [Dew00]. Conventional [Dew01c]. Conversions [Bec00a]. Converting [Bia00]. Cookie [Pad01]. Cool [Bri01g]. Copying [Rog00a]. CORBA [May01]. Core [Käm01]. Correction [And02]. Could [All00g]. Counting [Bat00, KM02c]. cout [Ano01-33]. Creating [Ehl00, Fos00, Hic00]. Creation [Wu00]. Cross [Gui00b, Kem02b, Mor01]. Cross-Platform [Gui00b, Kem02b, Mor01]. Custom [Vla02]. Customize [Yod01]. Cyclic [Ber02].

D [Tel00]. Dance [Sch02e]. Data [Bra00, KM02e, Lin01, Thi00]. Data-Based [Bra00]. Debug [Nel01]. Debugging [Fox01, Gui00c, Vla02, Zac01]. Debunker [KB00c, KB00d, KB00b, KB00e, KB00a, KB01c, KB01b, KB01e, KB01a, KB01d, KB02]. Declarations [Mey01b]. Decryptor [Zol01]. Defining [Aus01b]. Delegate [Lip01]. Delegates [Smi02]. Dependencies [Ber02]. Depth [Phi00d]. Derivable [Cha01]. Design [Bat01, Bec00b, Bec00c, Bec02c, Laf00, Sut02a]. Destructing [Pet00]. Determination [Bra00]. Development [Cum02, Gui00b, KL02]. Deviation [Sch00h]. Dialog [Kem02b]. Dictionary [And02]. Dies [Sch00c]. Different [Aus02a]. Dimensional [Bav00b]. Directories [Nik00]. Directory [Mor00a]. Discontinuum [Sch02d]. Distinguishing [Mey01h]. Do [Bri01d]. Document [Mor00b]. Documenting [Peh01]. Documents [Bav00a]. Don’t [Sch02b]. Double [Phi00a]. Dragons [Gri01]. Driven [Bec02c, KL02]. Drowning [Bri01e]. Dynamic [Gon00].

Easier [Ano01a, Koe01, KM01a, KM01d, KM01b, KM01e, KM01c, KM02d, KM02e, KM02a, KM02b, KM02c]. Easy [Bat01, GF01, Gon00, Ano01-33]. Economical [Bat00]. Editor [All02a, All02g, All02b, All02d, All02c, All02f, All02e, Ano01c, Ano01d, Ano02b, Ano02c, Ano02d, Bri00b, Bri00c, Bri00d, Bri00e, Bri00f, Bri00g, Bri00h, Bri00i, Bri00j, Bri00k, Bri00l, Bri00m, Bri01e, Bri01d, Bri01h, Bri01f, Bri01i, Bri01g, Bri01j, Bri01a, Bri01b, Bri01c]. Editors [Cir01]. Effective [Mey01j, Wu00]. Efficient [Lem02, Lor01, Pyr02]. Elegant [Hub01b]. Embedded [Fau00, Ano01e]. Emperor [Bri01h]. Encapsulating [May01, Rog02]. Encapsulation [Mey00d]. Engineering [Käm01]. Environment [Hub02]. Error [Bec00d, Off00, Zol01]. Errors [Vla02]. Euclid [All02b]. Events [Smi02]. eWriter [Sch01d]. Excel [LK00]. Exception [Ano01e, KM02b, KM02c, Nie02, Sta01, Sut02d]. Exceptional [Bri00a]. Exceptions [Nie02, Sch00a, Sch00k, Sch00b, Sch00c, Sch00d, Sch00e, Sch00h, Sch00f, Sch00g, Sch00i, Sch00j, Sch01c, Sch01b, Sch01a, Sch01d, Sch01e, Sch01i, Sch01f, Sch01g, Sch01h, Sch01j, Sch02b, Sch02c, Sch02d, Sch02e]. Exchanging [Lin01]. Execution [Nak01]. exit [Pla00c]. Experts [Sch01f]. Export [Sut02c]. Expressing [Wei02]. Expressions [Ber00]. Extending [Ano02e, Til01]. Extensible [Sut02b]. Extensions [Bec00j, Lip02]. Extracting [Thi00].

Factories [Ehl00]. Failures [Sut00]. Farm [Sut01c]. Fi [Sch00f]. File [Aus01c, Yor01]. File-Based [Aus01c]. Files [Rog00a]. Film [Sch01c]. Filter [Peh01]. Filtering [Kue05]. Filters [Phi00c]. Finding [Bas00]. First [KB00e]. Flexible [Lem02, Mey02c]. Flies [KB01c]. Floating [Bec00d, Bec00e, Bec00f]. Floating-Point [Bec00d, Bec00e, Bec00f]. Form [PW00]. Formatted [All00d]. Formatters [Sut01c]. FORTRAN [Mey00c]. Forum [All02a, All02g, All02b, All02d, All02c, All02f, All02e, Ano01c, Ano01d, Ano02b, Ano02c, Ano02d, Bri00b, Bri00c, Bri00d, Bri00e, Bri00f, Bri00g, Bri00h, Bri00i, Bri00j, Bri00k, Bri00l, Bri00m, Bri01e, Bri01d, Bri01h, Bri01f, Bri01i, Bri01g, Bri01j, Bri01a, Bri01b, Bri01c]. Framework [All00g, BMG02, Hsi00, Mor01]. Fugit [Sch00j]. Function [Aus01d, Bec02e, Dew01e, Hub01b, Nas00, Sut01f]. Functions [Mey00d, Mey02b, Yod01]. Fungible [Dew00]. Fuzzy [Sch01e].

Game [Bri01i, Cum02]. GCC [Gui00b]. Gear [All02c]. Gear-Head [All02c]. Geek [Sch00g]. General [BMG02]. Generalizing [Vla01]. Generation [Geo02]. Generic [Bav00b, Bec00j, Bec00m, Bec00l, Bec00k, Bec01c, Bec01b, Bec01f, Bec01e, Bec01d, Bec02c, Bec02d, Bec02e, Bec02a, Bec02b, Cac00a, Dew01c]. Genetic [Buc02]. Geometry [Abl02]. Global [Ten02]. GNOME [Pin02]. GNU [Col02, Zac01]. GNU/Linux [Zac01]. Going [Bri01f]. Good [All02d, Bec00h, Sch01b]. Graphic [Peh01]. Great [Sch01i]. Grow [KM01d]. Gtk [Ano02p]. GUI [Mor01]. Guidelines [Mey01j, Pee01].

habits [Sut02e]. Hand [Mey01i]. Hand-Written [Mey01i]. Handle [KM02b]. Handlers [Bec00f]. Handles [KM02b, KM02c]. Handling [Ano01e]. Handy [Fox01]. Happen [Bec00h]. Head [All02c]. Ḧello [Ano01-33]. Here [Gri01]. Homework [Bri01d]. HTML [Bav00a, Mor00b]. Humor [All02c].

I/O [All00b, Aus01d]. Identifying [Ber02]. IDEs [Sor01]. Idioms [Dew01a]. II [Wal01]. Image [Lor01, Phi00c, Qur01]. Images [Phi00d]. Implementation [Bar00]. Implementing [Gat02, Liu01]. Import [Peh01, All00a, All00b, All00c, All00d, All00e, All00f]. Important [All02c, All02e]. Improve [Mey00d]. Improved [Cac00b, Wei02]. Index [Fos00, Hic00]. Inheritance [Sut02b]. Initializations [Mey01b]. Inline [Mey02b]. Inner [All00e]. Integers [Mey00a, Mey01c, Mey01d]. Integrating [Cal00]. Integration [GF01]. Interface [Peh01]. Interfaces [All00e]. International [Sch01f]. Interpolation [Lou00]. Interrupt [Gat02]. Interval [Dan00]. Introducing [Mey00b]. Introduction [Bec00k, Lou00, Nas00]. Intrusive [KM02c]. Intuitive [Mor01]. Iostream [Ano01f]. Iostream-Compatible [Ano01f]. IOStreams [But00, Aus00, Off00]. Irae [Sch00c]. Items [All02c]. Iterator [Dew02d, Mey01j]. Iterators [Aus01b, Aus01f, Bec01b, Bec01f, Bec01e].

Java [Lin01, Pla00d]. java.*

[All00a, All00b, All00c, All00d, All00e, All00f]. JNI [GF01]. Journal [Ano01b]. Journeyman [Bec00b, Bec00c, Bec00a, Bec00d, Bec00e, Bec00f, Bec00g, Bec00i, Bec00h, Bec01a]. Just [Pla00a].

KDE [Mot02]. Kelly [Sch00k]. Keyword [And02]. Knowledge [Dew00, Dew01c, Dew01d, Dew01e, Dew01b, Dew01a, Dew01f, Dew02c, Dew02d, Dew02e].

Language [Pla02]. Languages [Sch01h]. Layer [Yor01]. LDAP [Mor00a]. Leaks [Gar00, Gui00c]. Length [Mey01e, Mey01f, Mey02d]. Librarian [Aus00, Aus01a, Aus01b, Aus01c, Aus01d, Aus01e, Aus01f, Aus02b, Aus02a]. Library [Aus01e, For01, Kem02a, Koe01, Pla00e, Rin01, Sut02f, Til01, Yod01, Sut02e, Cav02]. Licensing [Sca00]. Lightweight [Han00]. Linear [Swa00]. Linked [Pla00b]. Linux [Bon01, Hub02, Nel00, Sor01]. Linux/Unix [Sor01]. List [Pla00b]. Literals [Mey01a]. Loading [Gon00]. Locales [All00d, But00]. Location [Sme00]. Locks [Abr02]. Logging [Off00, Pyr02]. Logically [Dew02e]. Look [Ano01a, Ano01g, Sut02d]. Loops [Mey01i].

Machine [Laf00]. Macro [Nak01, Tim01]. Macros [Get00, Mey01g]. Made [Ano01a, Bat01, GF01, Koe01, KM01a, KM01d, KM01b, KM01e, KM01c, KM02d, KM02e, KM02a, KM02b, KM02c]. Mail [Ano00m, Ano00n, Ano00o, Ano00p, Ano00q, Ano00r, Ano00s, Ano00t, Ano00u, Ano00v, Ano00w, Ano00x, Ano01u, Ano01v, Ano01w, Ano01x, Ano01y, Ano01z, Ano01-27, Ano01-28, Ano01-29, Ano01-30, Ano01-31, Ano01-32]. Maintainable [Ehl00]. Making [Cha01, Low00]. MAM [KB00b]. Man [Sch01a]. Managed [Cav02, Lip02]. Manager [Pad01]. Managing [Ber02]. Manor [Sut01c]. MAPM [Rin01]. Mapping [Wu00]. Maps [Qur01]. Math [Rin01, Sch01e]. Mathematical [Liu01]. MAW [KB00b]. mC [Wal01]. mC/OS [Wal01]. mC/OS-II [Wal01]. Me [KB01d]. Meaning [Wei02]. Member [Cac00b, Mey00d]. Members [Mey02c]. Memory [Fau00, Gar00, Gui00c, Rog02, Sut01b, Vla02]. Memory-Based [Fau00]. Message [Zol01]. Metaphone [Phi00a]. Metaprogrammed [Dew02c]. Metaprogramming [Bec02b]. migrating [Sut02e]. Mill [Ano01t, Sut00, Sut01b, Sut01d, Sut01c, Sut01e, Sut01f, Sut02e, Sut02c, Sut02d, Sut02f]. Mistakes [Bec00b, Bec00c]. Model [Bec00l]. Modern [Sut02a]. Monitoring [Pyr02]. Moore [Ber00]. Mortem [KB00c, KB00d, KB00b, KB00e, KB00a, KB01c, KB01b, KB01e, KB01a, KB01d, KB02]. Multiple [Abr02, HR02, Jär01]. Multithreaded [Ten02]. Multithreading [Bec00g]. Mystery [Sch01f].

Name [Bri01i, Sch01h, Sch02a]. Naming [KM02d]. Navigating [Bon01]. Nearest [And01]. Neat [Bas00]. Ned [Sch00k]. Neighbor [And01]. Network [Nel00, Rog00b]. Neural [Rog00b]. Never [Sut00]. Nevermind [Sch01g]. Next [KB01e]. next permutation [Ano02a]. No [Sch00d]. Noise [Phi00c]. Non [Cha01, Mey00d]. Non-Derivable [Cha01]. Non-Member [Mey00d]. Nonlinear [Per00a, Per00b]. Not-So-Obvious [Get00]. Numbers [Aus02b, Bec00h].

O [All00b, Aus01d]. Object [Dix01].

Objects [Aus01d, Bec02e, Dew01e, Kou00, LK00, Lop01, Low00, Sut00, Ten02]. Observer [Bar00, Bat01]. Observer/ Mediator [Bat01]. Obvious [Get00]. Old [KM02e]. One [Dew01d]. Open [Sor01]. OpenGL [Qur01, Wu00]. Operations [Nie00]. Operator [Gib00, Dew02a, Dew02b, Sch02a]. Optimization [Buc01]. Optimizing [Nie00]. Organizing [Rog00b]. OS-II [Wal01]. Other [All02c]. Output [Dew02d]. Overruns [KM02a]. Own [Bec01e].

Palm [Har02]. Panjandrum [Sch01i].

Parameter [Que00]. Parameters [Mey02d]. Part [Ano01t, Bec00b, Bec00c, Bec01a, Dew02a, Dew02b, KM02b, KM02c, Mey00a, Mey01c, Mey01d, Mey01e, Mey02d, Per00a, Per00b, Sak00a, Sak00b, Sut01d, Sut02c, Sut02f, Mey02c]. Pattern [Bar00]. Patterns [Bat01, May01]. PC [Ano02e]. C [KL02, Pla02]. IP [Sme00]. Linux [Zac01]. Mediator [Bat01]. Unix [Sor01]. Write [Abr02]. Per-Thread [Cha02]. Perchance [Ano01t, Sut01d]. Performing [Abr02].

Perl [Sew02]. Persistence [Hsi00]. Perspective [Tel00]. Physical [Ber02]. Plain [KM02e]. Platform [Gui00b, Kem02b, Mor01]. Platforms [Lop01]. Please [Dew01d]. Plots [Thi00]. Plotting [Bas00]. Pocket [Ano02e]. Point [Bec00d, Bec00e, Bec00f]. Pointers [Kar02, Mey02d]. Policy [Bec02c, Bro02]. Policy-Driven [Bec02c]. Polymorphic [Dew01e]. Portability [All02e]. Portable [Gib00, Rin01]. Possibly [All00g]. Post [KB00c, KB00d, KB00b, KB00e, KB00a, KB01c, KB01b, KB01e, KB01a, KB01d, KB02]. Post-Mortem [KB00c, KB00d, KB00b, KB00e, KB00a, KB01c, KB01b, KB01e, KB01a, KB01d, KB02]. Practice [Sak00a, Sak00b]. Pragmatic [Phi00b, Sut02d]. Precision [KB02, Rin01]. Predicate [How00]. Primitive [Low00]. Probing [Har02]. Problem [And01]. Problems [Lab00]. Procedures [Liu01]. Products [Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano01m, Ano01n, Ano01o, Ano01p, Ano01q, Ano01r, Ano01s, Ano02f, Ano02g, Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano02m, Ano02n, Ano02o]. Professional [Hub01a]. Programmer [Wal01, Phi00b]. Programmers [Mey02a]. Programming [Bec00m, Bec00l, Bec00k, Bec01c, Bec01b, Bec01f, Bec01e, Bec01d, Bec02c, Bec02d, Bec02e, Bec02a, Bec02b, Cac00a, Kem02b, LK00, Nel00, Pin02, Pla02, Sme01]. Programs [Ano01e, KM01a, Lin01]. progress [Sut02e]. Prompt [Rog00a]. Protocol [Dix01]. Pure [KB00c]. Purpose [BMG02].

Question [Dew01b]. Quicksort [Bec01a].

Random [Yor01]. rarely [Sch02a]. Read [Abr02, Sch02b]. Read/Write [Abr02]. Reading [Que00]. Real [Pyr02, Wal01]. Real-Time [Pyr02, Wal01]. Recursion [Cre00]. Redmond [Sch02a]. Refactoring [Ten02]. Reference [Bat00, KM02c, For01]. Reference-Counting [Bat00]. Regression [Swa00]. Regular [Ber00]. Release [Bro00]. Remembrance [Sch00e]. remove if [Now01]. Removing [Phi00c]. Rendering [Phi00d]. Replacing [Sak00a, Sak00b]. Reply [KB01b]. Resizable [Bav00b]. Resources [Bia00]. Respect [Dew01b]. Restrictions [Sut02c]. Retargeting [Col02]. Retreat [Sch02c]. Retro [Bec00i]. Return [Gui00a, Jär01]. Reusable [Liu01, Per00a, Per00b]. Reuse [Hub02]. Review [Sut02a, Bri00a, But00, Cir01, For01, Gla01, Käm01, Phi00b, Wal01]. right [Sch02a]. Rotation [Qur01]. Round [Dew02e]. Roundup [Sut01a]. Routines [Gat02]. Rule [KM01b]. Running [Dew02e].

Safe [Bat00, Kle00, Low00]. Safety [KM02b, KM02c, Wei02]. scaffolding [Ano01-33]. Scales [Bas00]. Scanning [Nik00]. Search [HR02, Ber00, Mey01h, Phi00a]. Searches [KM01c]. Searching [Aus01e]. Secure [Sca00]. Self [Pet00, Rog00b]. Self-Organizing [Rog00b]. Semper [Sch00f]. Sending [Lop01]. Serialization [BMG02]. Server [Sme00]. Service [Gat02]. Set [Smi00]. Shall [KB00d]. Shim [Rei00]. Shoes [Bri01h]. Shop [Bec00b, Bec00c, Bec00a, Bec00d, Bec00e, Bec00f, Bec00g, Bec00i, Bec00h, Bec01a]. Should [KM01e]. Siblings [Str02c]. Simple [Bri00n, Hsi00, KM02b, Sme01, Swa00, Dix01]. Simplest [All00g]. Simulated [Buc01]. Singleton [Cha02]. Singly [Pla00b]. Smart [Kar02]. Smithee [Sch01c]. SOAP [Dix01]. Socket [Ano01f, Sme00, Sme01, Ano01-33]. Sockets [Pee01]. Soft [Sch00g]. Software [Hub01a, Käm01]. Solver [Per00a, Per00b]. Solving [Lab00]. Something [Aus02a, Bri01g]. Sorting [Bec01a]. Source [Bon01, Sor01]. Space [Sch02d]. Space-Time [Sch02d]. Specialize [Sut01f]. Specifications [Sut02d]. Spiral [Sch02e]. Split [Dew01a]. Spurious [KB02]. SQL [Fau00]. Stack [Nie02]. Standard [Aus01e, But00, Cav02, For01, Pla00d, Sch00h, Sch01h, Til01, Aus00, Aus01a, Aus01b, Aus01c, Aus01d, Aus01e, Aus01f, Aus02b, Aus02a, Pla00d, Pla00b, Pla00f, Pla00e, Pla00c, Sut02e, Sut02f]. State [Laf00]. Statements [Ano01a]. Static [Wei02]. Stdio [Aus00]. Stepping [Dew01f]. Sticky [Bec00f]. Still [All02e]. STL [Bec00j, Bec00m, Bec00l, Bec00k, Bec01c, Bec01b, Bec01f, Bec01e, Bec01d, Bec02c, Bec02d, Bec02e, Bec02a, Bec02b, Fos00, Hic00, Lab00, Mey01h, Mey01i, Smi00, Zol01]. STL-Based [Smi00]. Straight [Sch02a]. Stream [All00b, Kue05, Nel01, Sew02]. Streambuf [Aus01f]. Streambufs [Aus01f]. Stressing [Kou00]. String [Ber00, Nie00, Pan01, Sut01c]. Strings [All00f, Sak00a, Sak00b]. Structures [Dew00]. Studies [Str02b]. Stuff [All02d]. Substring [HR02, Nie00]. suitable [Ano01-33]. Sunshine [Sch00i]. Superman [Sch00i]. Sutter [Ano01t, Sut00, Sut01b, Sut01d, Sut01c, Sut01e, Sut01f, Sut02e, Sut02c, Sut02d, Sut02f]. System [Per00a, Per00b]. Systems [Pyr02].

T [Now01]. Table [Fos00, Hic00]. Take [Fos00]. talk [Sch02a]. Taming [Tel00]. TCP [Sme00]. TCP/IP [Sme00]. Technique [Lem02, Pyr02]. Teh [Aus01f]. Template [And01, Bav00b, Bec00l, Bec02a, Cal00, Til01, Yod01]. Templates [Cac00b, Geo02, Mor00b, Sut01f, Sut02b]. Tempus [Sch00j]. Test [All00g, KL02]. Testing [Pla00f, Pla00e, Sta01]. Tests [Bri00n, Sta01]. Text [All00d]. Texture [Qur01, Wu00]. Thee [KB00d]. Their [Bec02e]. Them [Sch01h]. Theory [Sak00a, Sak00b]. Things [Bec00h, Sch01b]. Thou [All02a]. Thread [Cha02, Kle00, Low00]. Thread-Safe [Kle00]. Threads [Cal00, How00, Pet00]. Three [KM01b, Mey01j]. Throw [Ano01t, Sut01d]. Time [Dew01d, KB01c, Pla00d, Pyr02, Sch02d, Wal01]. Tip [Abr02, Cac00a, Cha01, Fos00, Hic00, Now01, Pet00, Tim01]. Tips [Abr02, Cac00a, Cha01, Fos00, Hic00, Now01, Pet00, Tim01]. Title [Sch02b]. Toolbox [Pan01]. Toolkit [Hub01a]. Tracing [Bro00, Lem02, Nie02]. Tracking [Gar00]. Traits [Bec01d, Sut02b]. Transform [Tel00]. Trap [Bec00f]. Traversal [Cre00]. Tree [Cre00]. Trees [Dan00]. Truly [Ehl00]. Trust [KB01d]. try [Nas00]. Tuple [Jär01]. Tutorial [For01]. Two [Ano01g, Sme01]. Type [Cac00b, Wei02]. typedefs [Mey02c]. typeof [Gib00, Dew02a, Dew02b]. Types [Cav02, Jär01, KM02d, Mor02].

Uncaught

[Sch00a, Sch00k, Sch00b, Sch00c, Sch00d, Sch00e, Sch00h, Sch00f, Sch00g, Sch00i, Sch00j, Sch01c, Sch01b, Sch01a, Sch01d, Sch01e, Sch01i, Sch01f, Sch01g, Sch01h, Sch01j, Sch02b, Sch02c, Sch02d, Sch02e]. Understanding [KM01c]. Unicode [Kue05]. Unions [Mor02]. Unit [All00g, Bri00n, Sta01]. Unknown [KM02d]. Unraveling [Bec00g]. Untested [Gui00a]. Usage [Mey01j]. Use [KM01e]. Users [Ano01b]. Using [Buc02, Cav02, Geo02, Gui00b, How00, Koe01, Liu01, Mor02, Qur01, Smi02, Yod01]. Utility [Get00].

Values [Jär01]. Variable

[Mey01e, Mey01f, Mey02d]. Variant [Cac00b]. VB [Ano02e]. VCL [Bia00]. vector [Now01]. Vectors [Aus01a, KM01d].

Version [Pla00a]. Via [Sut02b, Phi00d]. Views [PW00]. Virtual [Rog02]. Virtuality [Sut01e]. Visual [Zol01]. Visualizing [Phi00d]. VLA [Mey02c]. VLAs [Mey02c]. vs [KB00c, Mey01i].

w [KM01d]. Waits [How00]. Warping [Lor01]. Watchman [Tru02]. Way [Smi00]. Web [Pad01, Sca00]. Web-Based [Sca00]. Welcome [Bec00m]. Well [Bri01f]. Were [Sut00]. Where [All02a]. Which [KM01e]. Whichever [KB00e]. Widgets [Ano02p]. Win32 [How00, Nel01]. Window [Han00]. Windows [Bia00, Gar00, Lin01, Rog02]. Wise [Dew02a, Dew02b]. without [Cre00]. Witness [Sch01j]. Woods [Bri01j]. Work [All00g, KM01a, Sut02e]. Working [Dix01]. world [Ano01-33]. Wrapper [Ano01f, Bro02, Han00]. Wrappers [Hub01b]. Wrapping [Pee01]. Writing [Bec01e, Mot02, Que00, Ano01-33]. Written [Ano02e, Mey01i]. wxWindows [Mor01].

X [Mey01g]. XML [Geo02]. XParam [BMG02]. XSL [Geo02].

References

Ablavsky:2002:ABC

[Abl02]  Vitaly Ablavsky.  Applying

BGL to computational geometry. C/C++ Users Journal, 20 (8):6, 8–12, August 2002. CODEN CCUJEX. ISSN 1075-2838.

Abramson:2002:CCT

[Abr02] Tomer Abramson. C/C++ tips:

Tip #8: Performing multiple read/write locks. C/C++ Users Journal, 20(5):25–??, May 2002. CODEN CCUJEX. ISSN 1075-

Allison:2000:IJA

[All00a] Chuck Allison. import java.*: Arrays. C/C++ Users Journal, 18 (3):46–??, March 2000. CODEN CCUJEX. ISSN 1075-2838.

Allison:2000:IJB

[All00b] Chuck Allison. import java.*: Basic stream I/O. C/C++ Users Journal, 18(11):58–??, November 2000. CODEN CCUJEX. ISSN 1075-2838.

Allison:2000:IJC

[All00c] Chuck Allison. import java.*: Collections and algorithms. C/C++ Users Journal, 18(9):76–??, September 2000. CODEN CCUJEX. ISSN 1075-2838.

Allison:2000:IJF

[All00d] Chuck Allison. import java.*: Formatted text and locales. C/C++ Users Journal, 18(7):60– ??, July 2000. CODEN CCUJEX. ISSN 1075-2838.

Allison:2000:IJI

[All00e] Chuck Allison. import java.*: Interfaces and inner classes. C/C++ Users Journal, 18(1):??, January 2000. CODEN CCUJEX. ISSN 1075-2838.

Allison:2000:IJS

[All00f] Chuck Allison. import java.*: Strings. C/C++ Users Journal, 18(6):??, June 2000. CODEN CCUJEX. ISSN 1075-2838.

Allison:2000:SAU

[All00g] Chuck Allison. The simplest automated unit test framework that could possibly work.

C/C++ Users Journal, 18(9):48– ??, September 2000. CODEN CCUJEX. ISSN 1075-2838.

Allison:2002:EFBa

[All02a] Chuck Allison. Editor’s forum:

Brother, where art thou? C/C++ Users Journal, 20(7):4, July 2002. CODEN CCUJEX. ISSN 1075-2838.

Allison:2002:EFE

[All02b] Chuck Allison. Editor’s forum:

Euclid alone. C/C++ Users Journal, 20(10):4, October 2002. CO-DEN CCUJEX. ISSN 1075-2838.

Allison:2002:EFGb

[All02c] Chuck Allison. Editor’s forum:

Gear-head humor and other important items. C/C++ Users Journal, 20(6):4, June 2002. CO-DEN CCUJEX. ISSN 1075-2838. Stan Kelly-Bootle announces his retirement from C/C++ Users Journal column writing.

Allison:2002:EFGa

[All02d] Chuck Allison. Editor’s forum:

Good stuff. C/C++ Users Journal, 20(4):6, April 2002. CODEN CCUJEX. ISSN 1075-2838.

Allison:2002:EFP

[All02e] Chuck Allison. Editor’s forum: Is portability still important? C/C++ Users Journal, 20 (5):4, May 2002. CODEN CCU-JEX. ISSN 1075-2838.

Allison:2002:EFN

[All02f] Chuck Allison. Editor’s forum:

New cheese. C/C++ Users Journal, 20(8):4, August 2002. CO-DEN CCUJEX. ISSN 1075-2838.

Allison:2002:EFBb

[All02g] Chuck Allison. Editor’s forum: The best of both! C/C++ Users Journal, 20(9):4, September 2002. CODEN CCUJEX. ISSN 1075-2838.

Andrews:2001:TNN

[And01] Larry Andrews. A template for the nearest neighbor problem. C/C++ Users Journal, 19(11):??, November 2001. CODEN CCU-JEX. ISSN 1075-2838.

Andrews:2002:KCD

[And02] Larry Andrews. Keyword correction from a dictionary. C/C++ Users Journal, 20(6):14–??, June 2002. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2000:NPa

[Ano00a] Anonymous. New products. C/C++ Users Journal, 18(1):??, January 2000. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2000:NPb

[Ano00b] Anonymous. New products. C/C++ Users Journal, 18(2):94– ??, February 2000. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2000:NPc

[Ano00c] Anonymous. New products. C/C++ Users Journal, 18(3):95– ??, March 2000. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2000:NPd

[Ano00d] Anonymous. New products. C/C++ Users Journal, 18(4):86– ??, April 2000. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2000:NPe

[Ano00e] Anonymous. New products.

C/C++ Users Journal, 18(5):80– ??, May 2000. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2000:NPf

[Ano00f] Anonymous. New products.

C/C++ Users Journal, 18(6):??, June 2000. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2000:NPg

[Ano00g] Anonymous. New products.

C/C++ Users Journal, 18(7):95– ??, July 2000. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2000:NPh

[Ano00h] Anonymous. New products.

C/C++ Users Journal, 18(8):87– ??, August 2000. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2000:NPi

[Ano00i] Anonymous. New products.

C/C++ Users Journal, 18(9): 103–??, September 2000. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2000:NPj

[Ano00j] Anonymous. New products.

C/C++ Users Journal, 18(10): 103–??, October 2000. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2000:NPk

[Ano00k] Anonymous. New products.

C/C++ Users Journal, 18(11): 88–??, November 2000. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2000:NPl

[Ano00l] Anonymous. New products. C/C++ Users Journal, 18(12): 80–??, December 2000. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2000:WMa

[Ano00m] Anonymous. We have mail. C/C++ Users Journal, 18(1):??, January 2000. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2000:WMb

[Ano00n] Anonymous. We have mail. C/C++ Users Journal, 18(2):98– ??, February 2000. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2000:WMc

[Ano00o] Anonymous. We have mail. C/C++ Users Journal, 18(3):99– ??, March 2000. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2000:WMd

[Ano00p] Anonymous. We have mail. C/C++ Users Journal, 18(4):90– ??, April 2000. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2000:WMe

[Ano00q] Anonymous. We have mail. C/C++ Users Journal, 18(5):84– ??, May 2000. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2000:WMf

[Ano00r] Anonymous. We have mail. C/C++ Users Journal, 18(6):??, June 2000. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2000:WMg

[Ano00s] Anonymous. We have mail.

C/C++ Users Journal, 18(7):99– ??, July 2000. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2000:WMh

[Ano00t] Anonymous. We have mail.

C/C++ Users Journal, 18(8):91– ??, August 2000. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2000:WMi

[Ano00u] Anonymous. We have mail.

C/C++ Users Journal, 18(9): 107–??, September 2000. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2000:WMj

[Ano00v] Anonymous. We have mail.

C/C++ Users Journal, 18(10): 107–??, October 2000. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2000:WMk

[Ano00w] Anonymous. We have mail.

C/C++ Users Journal, 18(11): 92–??, November 2000. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2000:WMl

[Ano00x] Anonymous. We have mail.

C/C++ Users Journal, 18(12): 84–??, December 2000. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2001:CME

[Ano01a] Anonymous. C++ made easier:

A new look at for statements. C/C++ Users Journal, 19(12):??, December 2001. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2001:CCU

[Ano01b] Anonymous. C/C++ Users Journal 20th anniversary. C/C++ Users Journal, 19(11):??, November 2001. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2001:EFk

[Ano01c] Anonymous. Editor’s forum. C/C++ Users Journal, 19(11):??, November 2001. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2001:EFl

[Ano01d] Anonymous. Editor’s forum. C/C++ Users Journal, 19(12):??, December 2001. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2001:EHE

[Ano01e] Anonymous. Exception handling in Embedded C programs. C/C++ Users Journal, 19(2):??, February 2001. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2001:ICS

[Ano01f] Anonymous. An Iostreamcompatible socket wrapper. C/C++ Users Journal, 19(12):??, December 2001. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2001:LTC

[Ano01g] Anonymous. A look at two C# books. C/C++ Users Journal, 19 (12):??, December 2001. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2001:NPa

[Ano01h] Anonymous. New products. C/C++ Users Journal, 19(1):88– ??, January 2001. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2001:NPb

[Ano01i] Anonymous. New products.

C/C++ Users Journal, 19(2):88– ??, February 2001. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2001:NPc

[Ano01j] Anonymous. New products.

C/C++ Users Journal, 19(3):88– ??, March 2001. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2001:NPd

[Ano01k] Anonymous. New products.

C/C++ Users Journal, 19(4):??, April 2001. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2001:NPe

[Ano01l] Anonymous. New products.

C/C++ Users Journal, 19(5):80– ??, May 2001. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2001:NPf

[Ano01m] Anonymous. New products.

C/C++ Users Journal, 19(6):71– ??, June 2001. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2001:NPg

[Ano01n] Anonymous. New products.

C/C++ Users Journal, 19(7):72– ??, July 2001. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2001:NPh

[Ano01o] Anonymous. New products.

C/C++ Users Journal, 19(8):72– ??, August 2001. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2001:NPi

[Ano01p] Anonymous. New products. C/C++ Users Journal, 19(9):64– ??, September 2001. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2001:NPj

[Ano01q] Anonymous. New products. C/C++ Users Journal, 19(10): 64–??, October 2001. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2001:NPk

[Ano01r] Anonymous. New products. C/C++ Users Journal, 19(11):??, November 2001. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2001:NPl

[Ano01s] Anonymous. New products. C/C++ Users Journal, 19(12):??, December 2001. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2001:SMN

[Ano01t] Anonymous. Sutter’s Mill: To new, perchance to throw (part 1 of 2). C/C++ Users Journal, 19 (3):67–??, March 2001. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2001:WMa

[Ano01u] Anonymous. We have mail. C/C++ Users Journal, 19(1):92– ??, January 2001. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2001:WMb

[Ano01v] Anonymous. We have mail. C/C++ Users Journal, 19(2):92– ??, February 2001. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2001:WMc

[Ano01w] Anonymous. We have mail.

C/C++ Users Journal, 19(3):92– ??, March 2001. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2001:WMd

[Ano01x] Anonymous. We have mail.

C/C++ Users Journal, 19(4):??, April 2001. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2001:WMe

[Ano01y] Anonymous. We have mail.

C/C++ Users Journal, 19(5):84– ??, May 2001. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2001:WMf

[Ano01z] Anonymous. We have mail.

C/C++ Users Journal, 19(6):75– ??, June 2001. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2001:WMg

[Ano01-27] Anonymous. We have mail.

C/C++ Users Journal, 19(7):76– ??, July 2001. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2001:WMh

[Ano01-28] Anonymous. We have mail.

C/C++ Users Journal, 19(8):76– ??, August 2001. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2001:WMi

[Ano01-29] Anonymous. We have mail.

C/C++ Users Journal, 19(9):68– ??, September 2001. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2001:WMj

[Ano01-30] Anonymous. We have mail. C/C++ Users Journal, 19(10): 68–??, October 2001. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2001:WMk

[Ano01-31] Anonymous. We have mail. C/C++ Users Journal, 19(11):??, November 2001. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2001:WMl

[Ano01-32] Anonymous. We have mail. C/C++ Users Journal, 19(12):??, December 2001. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2001:SSW

[Ano01-33] Anonymous. With suitable scaffolding, writing to a socket is

as easy as cout << "Hello,

world". C/C++ Users Journal, 19(12):??, December 2001. CO-DEN CCUJEX. ISSN 1075-2838.

Anonymous:2002:CAN

[Ano02a] Anonymous. C++ algorithms:

next permutation.  C/C++

Users Journal, 20(3):??, March 2002. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2002:EFa

[Ano02b] Anonymous. Editor’s forum. C/C++ Users Journal, 20(1):??, January 2002. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2002:EFb

[Ano02c] Anonymous. Editor’s forum. C/C++ Users Journal, 20(2):??, February 2002. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2002:EFc

[Ano02d] Anonymous. Editor’s forum.

C/C++ Users Journal, 20(3):??, March 2002. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2002:EPP

[Ano02e] Anonymous. Extending Pocket PC apps written in C with VB. C/C++ Users Journal, 20(1):??, January 2002. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2002:NPa

[Ano02f] Anonymous. New products.

C/C++ Users Journal, 20(1):??, January 2002. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2002:NPb

[Ano02g] Anonymous. New products.

C/C++ Users Journal, 20(2):??, February 2002. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2002:NPc

[Ano02h] Anonymous. New products.

C/C++ Users Journal, 20(3):??, March 2002. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2002:NPd

[Ano02i] Anonymous. New products.

C/C++ Users Journal, 20(4):65– ??, April 2002. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2002:NPe

[Ano02j] Anonymous. New products.

C/C++ Users Journal, 20(5):58– ??, May 2002. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2002:NPf

[Ano02k] Anonymous. New products. C/C++ Users Journal, 20(6):58– ??, June 2002. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2002:NPg

[Ano02l] Anonymous. New products. C/C++ Users Journal, 20(7):65– 67, July 2002. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2002:NPh

[Ano02m] Anonymous. New products. C/C++ Users Journal, 20(8):57– 59, August 2002. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:2002:NPi

[Ano02n] Anonymous. New products. C/C++ Users Journal, 20(9):57– 59, September 2002. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2002:NPj

[Ano02o] Anonymous. New products. C/C++ Users Journal, 20(10): 57–59, October 2002. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2002:WG

[Ano02p] Anonymous. Widgets in gtk+. C/C++ Users Journal, 20(3):??, March 2002. CODEN CCUJEX. ISSN 1075-2838.

Austern:2000:SLI

[Aus00] Matt Austern. The Standard Librarian: IOStreams and Stdio. C/C++ Users Journal, 18(11): 55–??, November 2000. CODEN CCUJEX. ISSN 1075-2838.

Austern:2001:SLB

[Aus01a] Matt Austern. The Standard Librarian: Bitsets and bit vectors. C/C++ Users Journal, 19(5):58– ??, May 2001. CODEN CCUJEX. ISSN 1075-2838.

Austern:2001:SLD

[Aus01b] Matt Austern. The Standard Librarian: Defining iterators and const iterators. C/C++ Users Journal, 19(1):74–??, January 2001. CODEN CCUJEX. ISSN 1075-2838.

Austern:2001:SLFa

[Aus01c] Matt Austern. The Standard Librarian: File-based containers. C/C++ Users Journal, 19(7):54– ??, July 2001. CODEN CCUJEX. ISSN 1075-2838.

Austern:2001:SLFb

[Aus01d] Matt Austern. The Standard Librarian: I/O and function objects. C/C++ Users Journal, 19(9):??, September 2001. CODEN CCU-JEX. ISSN 1075-2838.

Austern:2001:SLS

[Aus01e] Matt Austern. The Standard Librarian: Searching in the standard library. C/C++ Users Journal, 19 (11):??, November 2001. CODEN CCUJEX. ISSN 1075-2838.

Austern:2001:TSL

[Aus01f] Matt Austern. Teh Standard Librarian: Streambufs and Streambuf iterators. C/C++ Users Journal, 19(3):72–??, March 2001. CO-DEN CCUJEX. ISSN 1075-2838.

Austern:2002:SLN

[Aus02a] Matt Austern. The Standard Librarian: And now for something completely different. C/C++ Users Journal, 20(1):??, January 2002. CODEN CCUJEX. ISSN 1075-2838.

Austern:2002:SLC

[Aus02b] Matt Austern. The Standard Librarian: Complex numbers. C/C++ Users Journal, 20(3):??, March 2002. CODEN CCUJEX. ISSN 1075-2838.

Barnes:2000:IOP

[Bar00] Paul Barnes. Implementation of observer pattern. C/C++ Users Journal, 18(7):50–??, July 2000. CODEN CCUJEX. ISSN 1075-2838.

Bas:2000:FNS

[Bas00] Antonio Gómiz Bas. Finding neat scales for plotting. C/C++ Users Journal, 18(3):42–??, March 2000. CODEN CCUJEX. ISSN 1075-2838.

Batov:2000:SER

[Bat00] Vladimir Batov. Safe and economical reference-counting in C++. C/C++ Users Journal, 18(6):??, June 2000. CODEN CCUJEX. ISSN 1075-2838.

Batov:2001:CME

[Bat01] Vladimir Batov. Callbacks made easy with the observer/mediator design patterns. C/C++ Users Journal, 19(2):??, February 2001. CODEN CCUJEX. ISSN 1075-2838.

Bavestrelli:2000:BHD

[Bav00a] Giovanni Bavestrelli. Building HTML documents with C++. C/C++ Users Journal, 18(4):67– ??, April 2000. CODEN CCU-JEX. ISSN 1075-2838.

Bavestrelli:2000:CTD

[Bav00b] Giovanni Bavestrelli. A class template for N-dimensional generic resizable arrays. C/C++ Users Journal, 18(12):32–??, December 2000. CODEN CCUJEX. ISSN 1075-2838.

Becker:2000:JSCc

[Bec00a] Pete Becker. The journeyman’s shop: Casts and conversions. C/C++ Users Journal, 18(4):73– ??, April 2000. CODEN CCU-JEX. ISSN 1075-2838.

Becker:2000:JSCa

[Bec00b] Pete Becker. The journeyman’s shop: Common design mistakes, part 1. C/C++ Users Journal, 18(1):??, January 2000. CODEN CCUJEX. ISSN 1075-2838.

Becker:2000:JSCb

[Bec00c] Pete Becker. The journeyman’s shop: Common design mistakes, part 2. C/C++ Users Journal, 18(2):77–??, February 2000. CO-DEN CCUJEX. ISSN 1075-2838.

Becker:2000:JSE

[Bec00d] Pete Becker. The journeyman’s shop: Error in floating-point calculations. C/C++ Users Journal, 18(7):73–??, July 2000. CODEN CCUJEX. ISSN 1075-2838.

Becker:2000:JSF

[Bec00e] Pete Becker. The journeyman’s shop: Floating-point basics. C/C++ Users Journal, 18(6):??, June 2000. CODEN CCUJEX. ISSN 1075-2838.

Becker:2000:JST

[Bec00f] Pete Becker. The journeyman’s shop: Trap handlers, sticky bits, and floating-point comparisons. C/C++ Users Journal, 18(12): 54–??, December 2000. CODEN CCUJEX. ISSN 1075-2838.

Becker:2000:JSU

[Bec00g] Pete Becker. The journeyman’s shop: Unraveling multithreading. C/C++ Users Journal, 18(8):71– ??, August 2000. CODEN CCU-JEX. ISSN 1075-2838.

Becker:2000:JSWb

[Bec00h] Pete Becker. The journeyman’s shop: When bad things happen to good numbers. C/C++ Users Journal, 18(10):54–??, October 2000. CODEN CCUJEX. ISSN 1075-2838.

Becker:2000:JSWa

[Bec00i] Pete Becker. The journeyman’s shop: When “retro” is better. C/C++ Users Journal, 18(5):67– ??, May 2000. CODEN CCUJEX. ISSN 1075-2838.

Becker:2000:GES

[Bec00j] Thomas Becker. Generic extensions to the STL. C/C++ Users Journal, 18(6):??, June 2000. CO-DEN CCUJEX. ISSN 1075-2838.

Becker:2000:SGPc

[Bec00k] Thomas Becker. STL & generic programming: Introduction to the STL. C/C++ Users Journal, 18 (12):68–??, December 2000. CO-DEN CCUJEX. ISSN 1075-2838.

Becker:2000:SGPb

[Bec00l] Thomas Becker. STL & generic programming: The template compilation model. C/C++ Users Journal, 18(10):65–??, October 2000. CODEN CCUJEX. ISSN 1075-2838.

Becker:2000:SGPa

[Bec00m] Thomas Becker. STL & generic programming: Welcome to STL and generic programming. C/C++ Users Journal, 18(9):70– ??, September 2000. CODEN CCUJEX. ISSN 1075-2838.

Becker:2001:JSS

[Bec01a] Pete Becker. The journeyman’s shop: Sorting through Quicksort, part 1. C/C++ Users Journal, 19 (1):66–??, January 2001. CODEN CCUJEX. ISSN 1075-2838.

Becker:2001:SGPb

[Bec01b] Thomas Becker. STL & generic programming: STL container iterators. C/C++ Users Journal, 19 (4):??, April 2001. CODEN CCU-JEX. ISSN 1075-2838.

Becker:2001:SGPa

[Bec01c] Thomas Becker. STL & generic programming: STL containers. C/C++ Users Journal, 19(2):??, February 2001. CODEN CCU-JEX. ISSN 1075-2838.

Becker:2001:SGPe

[Bec01d] Thomas Becker. STL & generic programming: Traits classes. C/C++ Users Journal, 19(12):??, December 2001. CODEN CCU-JEX. ISSN 1075-2838.

Becker:2001:SGPd

[Bec01e] Thomas Becker. STL & generic programming: Writing your own iterators. C/C++ Users Journal, 19(8):51–??, August 2001. CO-DEN CCUJEX. ISSN 1075-2838.

Becker:2001:SGPc

[Bec01f] Thomas Becker. STL and generic programming: More on iterators. C/C++ Users Journal, 19(6):38– ??, June 2001. CODEN CCUJEX. ISSN 1075-2838.

Becker:2002:SGPd

[Bec02a] Thomas Becker. STL & generic programming: C++ template programming. C/C++ Users Journal, 20(8):40, 42–46, August 2002. CODEN CCUJEX. ISSN 1075-2838.

Becker:2002:SGPe

[Bec02b] Thomas Becker. STL & generic programming: More on C++ metaprogramming. C/C++ Users Journal, 20(10):38, 40–44, October 2002. CODEN CCUJEX. ISSN 1075-2838.

Becker:2002:SGPa

[Bec02c] Thomas Becker. STL & generic programming: Policy-driven design. C/C++ Users Journal, 20 (2):??, February 2002. CODEN CCUJEX. ISSN 1075-2838.

Becker:2002:SGPb

[Bec02d] Thomas Becker. STL & generic programming: STL algorithms. C/C++ Users Journal, 20(4):44– ??, April 2002. CODEN CCU-JEX. ISSN 1075-2838.

Becker:2002:SGPc

[Bec02e] Thomas Becker. STL & generic programming: STL function objects and their adaptors. C/C++ Users Journal, 20(6):37–??, June 2002. CODEN CCUJEX. ISSN 1075-2838.

Berry:2000:CBM

[Ber00] David Berry. Combining Boyer– Moore string search with regular expressions. C/C++ Users Journal, 18(6):??, June 2000. CODEN CCUJEX. ISSN 1075-2838.

Bergin:2002:IMC

[Ber02] Thomas Bergin. Identifying and managing cyclic physical dependencies. C/C++ Users Journal, 20(7):6, 8–10, 12–14, July 2002. CODEN CCUJEX. ISSN 1075-2838.

Bianchi:2000:CVC

[Bia00] Luigi Bianchi. Converting VCL components to Windows resources. C/C++ Users Journal, 18(8):10–??, August 2000. CO-DEN CCUJEX. ISSN 1075-2838.

Brand:2002:XGP

[BMG02] Michael Brand, Ronnie Maor, and Sasha Gontmakher. XParam — A general purpose serialization framework for C++. C/C++ Users Journal, 20(7):38, 40–42, 44–46, July 2002. CODEN CCU-JEX. ISSN 1075-2838.

Bonang:2001:NLS

[Bon01] James Bonang. Navigating Linux source code. C/C++ Users Journal, 19(3):18–??, March 2001. CO-DEN CCUJEX. ISSN 1075-2838.

Bramley:2000:DBA

[Bra00] Michael Bramley. Data-based axis determination. C/C++ Users Journal, 18(7):20–??, July 2000. CODEN CCUJEX. ISSN 1075-2838.

Briand:2000:BRE

[Bri00a] Marc Briand. Book review: Exceptional C++. C/C++ Users Journal, 18(2):73–??, February 2000. CODEN CCUJEX. ISSN 1075-2838.

Briand:2000:EFa

[Bri00b] Marc Briand. Editor’s forum. C/C++ Users Journal, 18(1):6, January 2000. CODEN CCUJEX. ISSN 1075-2838.

Briand:2000:EFb

[Bri00c] Marc Briand. Editor’s forum. C/C++ Users Journal, 18(2):6, February 2000. CODEN CCU-JEX. ISSN 1075-2838.

Briand:2000:EFc

[Bri00d] Marc Briand. Editor’s forum. C/C++ Users Journal, 18(3):6, March 2000. CODEN CCUJEX. ISSN 1075-2838.

Briand:2000:EFd

[Bri00e] Marc Briand. Editor’s forum. C/C++ Users Journal, 18(4):6, April 2000. CODEN CCUJEX. ISSN 1075-2838.

Briand:2000:EFe

[Bri00f] Marc Briand. Editor’s forum.

C/C++ Users Journal, 18(5):6, May 2000. CODEN CCUJEX. ISSN 1075-2838.

Briand:2000:EFf

[Bri00g] Marc Briand. Editor’s forum.

C/C++ Users Journal, 18(6):6, June 2000. CODEN CCUJEX. ISSN 1075-2838.

Briand:2000:EFg

[Bri00h] Marc Briand. Editor’s forum.

C/C++ Users Journal, 18(7):6, July 2000. CODEN CCUJEX. ISSN 1075-2838.

Briand:2000:EFh

[Bri00i] Marc Briand. Editor’s forum.

C/C++ Users Journal, 18(8):6, August 2000. CODEN CCUJEX. ISSN 1075-2838.

Briand:2000:EFi

[Bri00j] Marc Briand. Editor’s forum.

C/C++ Users Journal, 18(9):6, September 2000. CODEN CCU-JEX. ISSN 1075-2838.

Briand:2000:EFj

[Bri00k] Marc Briand. Editor’s forum.

C/C++ Users Journal, 18(10):6, October 2000. CODEN CCUJEX. ISSN 1075-2838.

Briand:2000:EFk

[Bri00l] Marc Briand. Editor’s forum.

C/C++ Users Journal, 18(11):6, November 2000. CODEN CCU-JEX. ISSN 1075-2838.

Briand:2000:EFl

[Bri00m] Marc Briand. Editor’s forum. C/C++ Users Journal, 18(12):6, December 2000. CODEN CCU-JEX. ISSN 1075-2838.

Briand:2000:SUT

[Bri00n] Marc Briand. Simple unit tests in C++. C/C++ Users Journal, 18(7):10–??, July 2000. CODEN CCUJEX. ISSN 1075-2838.

Briand:2001:EFb

[Bri01a] Marc Briand. Editor’s forum. C/C++ Users Journal, 19(2):6, February 2001. CODEN CCU-JEX. ISSN 1075-2838.

Briand:2001:EFd

[Bri01b] Marc Briand. Editor’s forum. C/C++ Users Journal, 19(4):6, April 2001. CODEN CCUJEX. ISSN 1075-2838.

Briand:2001:EFi

[Bri01c] Marc Briand. Editor’s forum. C/C++ Users Journal, 19(9):6, September 2001. CODEN CCU-JEX. ISSN 1075-2838.

Briand:2001:EFDb

[Bri01d] Marc Briand. Editor’s forum: Do your homework. C/C++ Users Journal, 19(10):6, October 2001. CODEN CCUJEX. ISSN 1075-2838.

Briand:2001:EFDa

[Bri01e] Marc Briand. Editor’s forum: Drowning big brother. C/C++ Users Journal, 19(3):6, March 2001. CODEN CCUJEX. ISSN 1075-2838.

Briand:2001:EFG

[Bri01f] Marc Briand. Editor’s forum: Going back to the well. C/C++ Users Journal, 19(6):6, June 2001. CODEN CCUJEX. ISSN 1075-2838.

Briand:2001:EFS

[Bri01g] Marc Briand. Editor’s forum:

Something cool. C/C++ Users Journal, 19(7):6, July 2001. CO-DEN CCUJEX. ISSN 1075-2838.

Briand:2001:EFE

[Bri01h] Marc Briand. Editor’s forum: The Emperor’s new shoes. C/C++ Users Journal, 19(5):6, May 2001. CODEN CCUJEX. ISSN 1075-2838.

Briand:2001:EFN

[Bri01i] Marc Briand. Editor’s forum: The name game. C/C++ Users Journal, 19(8):6, August 2001. CO-DEN CCUJEX. ISSN 1075-2838.

Briand:2001:EFW

[Bri01j] Marc Briand. Editor’s forum: The woods of complexity. C/C++ Users Journal, 19(1):6, January 2001. CODEN CCUJEX. ISSN 1075-2838.

Bromberg:2000:TAR

[Bro00] Boris Bromberg. Tracing an application in release configuration. C/C++ Users Journal, 18(9):62– ??, September 2000. CODEN CCUJEX. ISSN 1075-2838.

Brodski:2002:PW

[Bro02] Matvei Brodski. Policy wrapper. C/C++ Users Journal, 20 (10):29–31, October 2002. CO-DEN CCUJEX. ISSN 1075-2838.

Bucci:2001:OSA

[Buc01] Mark Bucci. Optimization with simulated annealing. C/C++ Users Journal, 19(11):??, November 2001. CODEN CCUJEX. ISSN 1075-2838.

Bucci:2002:UGA

[Buc02] Mark Bucci. Using genetic algorithms. C/C++ Users Journal, 20(6):20–??, June 2002. CODEN CCUJEX. ISSN 1075-2838.

Button:2000:BRS

[But00] Brian Button. Book review: Standard C++ IOStreams and Locales. C/C++ Users Journal, 18(11): 51–??, November 2000. CODEN CCUJEX. ISSN 1075-2838.

Cacciola:2000:CCT

[Cac00a] Fernando Cacciola. C/C++ tips: Tip #2: Generic constants for generic programming. C/C++ Users Journal, 18(9):65– ??, September 2000. CODEN CCUJEX. ISSN 1075-2838.

Cacciola:2000:IVT

[Cac00b] Fernando Cacciola. An improved variant type based on member templates. C/C++ Users Journal, 18(10):10–??, October 2000. CO-DEN CCUJEX. ISSN 1075-2838.

Calkins:2000:ITT

[Cal00]  Charles Calkins.  Integrating

threads with template classes. C/C++ Users Journal, 18(5):32– ??, May 2000. CODEN CCUJEX. ISSN 1075-2838.

Caves:2002:UCS

[Cav02] Jonathan Caves. Using the C++ Standard Library with managed types. C/C++ Users Journal, 20 (10):34, 36, October 2002. CO-DEN CCUJEX. ISSN 1075-2838.

Chandrabose:2001:CCT

[Cha01] Shanker Chandrabose. C/C++ tips: Tip #6: Making classes nonderivable. C/C++ Users Journal, 19(5):51–??, May 2001. CODEN CCUJEX. ISSN 1075-2838.

Chaudhry:2002:PTS

[Cha02] Puneesh Chaudhry. A per-thread singleton class. C/C++ Users Journal, 20(5):14–??, May 2002. CODEN CCUJEX. ISSN 1075-2838.

Cirillo:2001:BRC

[Cir01] William Cirillo. Book review:

C/C++ Contributing Editors. C/C++ Users Journal, 19(12):??, December 2001. CODEN CCU-JEX. ISSN 1075-2838.

Collison:2002:RGC

[Col02] Michael Collison. Retargeting the GNU C compiler. C/C++ Users Journal, 20(6):30, 32–36, June 2002. CODEN CCUJEX. ISSN 1075-2838.

Creux:2000:TTC

[Cre00] Valery Creux. Tree traversal in C without recursion. C/C++ Users Journal, 18(7):56–??, July 2000. CODEN CCUJEX. ISSN 1075-

Cummins:2002:DG

[Cum02] Darin Cummins. The development game. C/C++ Users Journal, 20(10):18, 20–21, October 2002. CODEN CCUJEX. ISSN 1075-2838.

Dandass:2000:IT

[Dan00] Yogi Dandass. Interval trees. C/C++ Users Journal, 18(1):??, January 2000. CODEN CCUJEX. ISSN 1075-2838.

Dewhurst:2000:CKF

[Dew00] Steve Dewhurst. Common knowledge: Fungible control structures. C/C++ Users Journal, 18(12): 46–??, December 2000. CODEN CCUJEX. ISSN 1075-2838.

Dewhurst:2001:CKSa

[Dew01a] Stephen Dewhurst. Common knowledge: Split idioms. C/C++ Users Journal, 19(6):49–??, June 2001. CODEN CCUJEX. ISSN 1075-2838.

Dewhurst:2001:CKQ

[Dew01b] Steve Dewhurst. Common knowledge: A question of respect. C/C++ Users Journal, 19(4):??, April 2001. CODEN CCUJEX. ISSN 1075-2838.

Dewhurst:2001:CKC

[Dew01c] Steve Dewhurst. Common knowledge: Conventional generic algorithms. C/C++ Users Journal, 19 (12):??, December 2001. CODEN CCUJEX. ISSN 1075-2838.

Dewhurst:2001:CKO

[Dew01d] Steve Dewhurst. Common knowledge: One at a time, please.

C/C++ Users Journal, 19(8):46– ??, August 2001. CODEN CCU-JEX. ISSN 1075-2838.

Dewhurst:2001:CKP

[Dew01e] Steve Dewhurst. Common knowledge: Polymorphic function objects. C/C++ Users Journal, 19 (2):??, February 2001. CODEN CCUJEX. ISSN 1075-2838.

Dewhurst:2001:CKSb

[Dew01f] Steve Dewhurst. Common knowledge: Stepping back. C/C++ Users Journal, 19(10):57–??, October 2001. CODEN CCUJEX. ISSN 1075-2838.

Dewhurst:2002:BWTa

[Dew02a] Steve Dewhurst. A bit-wise Typeof operator, part 1. C/C++ Users Journal, 20(8):48, 50–52, 54, August 2002. CODEN CCU-JEX. ISSN 1075-2838.

Dewhurst:2002:BWTb

[Dew02b] Steve Dewhurst. A bit-wise Typeof operator, part 2. C/C++ Users Journal, 20(10):51–54, October 2002. CODEN CCUJEX. ISSN 1075-2838.

Dewhurst:2002:CKM

[Dew02c] Steve Dewhurst. Common knowledge: Metaprogrammed adapters. C/C++ Users Journal, 20(4):56– ??, April 2002. CODEN CCU-JEX. ISSN 1075-2838.

Dewhurst:2002:CKO

[Dew02d] Steve Dewhurst. Common knowledge: Output iterator adapters. C/C++ Users Journal, 20(2):??, February 2002. CODEN CCU-JEX. ISSN 1075-2838.

Dewhurst:2002:CKR

[Dew02e] Steve Dewhurst. Common knowledge: Running circles round you, logically. C/C++ Users Journal, 20(6):51–??, June 2002. CODEN CCUJEX. ISSN 1075-2838.

Dix:2001:WSS

[Dix01] Chris Dix. Working with SOAP, the Simple Object Access Protocol. C/C++ Users Journal, 19 (1):22–??, January 2001. CODEN CCUJEX. ISSN 1075-2838.

Ehlinger:2000:CTM

[Ehl00] Early Ehlinger. Creating truly maintainable class factories. C/C++ Users Journal, 18(11):22–??, November 2000. CODEN CCU-JEX. ISSN 1075-2838.

Faul:2000:EMB

[Fau00] Ben Faul. An embedded memorybased SQL class for C++. C/C++ Users Journal, 18(3):64– ??, March 2000. CODEN CCU-JEX. ISSN 1075-2838.

Forshaw:2001:BRC

[For01] Ben Forshaw. Book review: The C++ Standard Library: A Tutorial and Reference. C/C++ Users Journal, 19(3):40–??, March 2001. CODEN CCUJEX. ISSN 1075-2838.

Foster:2000:CCT

[Fos00] Sol Foster. C/C++ tips: Tip #3: Creating an index table in STL, take 2. C/C++ Users Journal, 18 (11):41–??, November 2000. CO-DEN CCUJEX. ISSN 1075-2838.

Fox:2001:HDC

[Fox01] Maurice Fox. A handy debugging class. C/C++ Users Journal, 19 (4):??, April 2001. CODEN CCU-JEX. ISSN 1075-2838.

Gareau:2000:TML

[Gar00] Jean L. Gareau. Tracking memory leaks under Windows CE. C/C++ Users Journal, 18(2):36– ??, February 2000. CODEN CCU-JEX. ISSN 1075-2838.

Gatliff:2002:IIS

[Gat02] Bill Gatliff. Implementing interrupt service routines in C++. C/C++ Users Journal, 20(2):??, February 2002. CODEN CCU-JEX. ISSN 1075-2838.

Georgescu:2002:CGT

[Geo02] Cristian Georgescu. Code generation templates using XML and XSL. C/C++ Users Journal, 20 (1):??, January 2002. CODEN CCUJEX. ISSN 1075-2838.

Getov:2000:OUM

[Get00] Radoslav Getov. Not-so-obvious utility macros. C/C++ Users Journal, 18(3):27–??, March 2000. CODEN CCUJEX. ISSN 1075-2838.

Gabrilovich:2001:JCI

[GF01] Evgeniy Gabrilovich and Lev Finkelstein. JNI–C++ integration made easy. C/C++ Users Journal, 19(1):10–??, January 2001. CODEN CCUJEX. ISSN 1075-

Gibbons:2000:PTO

[Gib00]  Bill Gibbons.  A portable

“typeof” operator. C/C++ Users Journal, 18(11):34–??, November 2000. CODEN CCUJEX. ISSN 1075-2838.

Glassborow:2001:BRA

[Gla01] Francis Glassborow. Book review: Accelerated C++. C/C++ Users Journal, 19(6):30–??, June 2001. CODEN CCUJEX. ISSN 1075-2838.

Gontmakher:2000:EDL

[Gon00] Sasha Gontmakher. Easy dynamic loading in C++. C/C++ Users Journal, 18(1):??, January 2000. CODEN CCUJEX. ISSN 1075-2838.

Griffiths:2001:HD

[Gri01] Alan Griffiths. Here be dragons. C/C++ Users Journal, 19(3):58– ??, March 2001. CODEN CCU-JEX. ISSN 1075-2838.

Guillemot:2000:CUR

[Gui00a] Marc Guillemot. Catching untested return codes. C/C++ Users Journal, 18(5):54–??, May 2000. CODEN CCUJEX. ISSN 1075-2838.

Guinan:2000:CPD

[Gui00b] Jamie Guinan. Cross-platform development using GCC. C/C++ Users Journal, 18(3):18–??, March 2000. CODEN CCUJEX. ISSN 1075-2838.

Guisado:2000:DCB

[Gui00c] Ernesto Guisado. Debugging component-based memory leaks.

C/C++ Users Journal, 18(1):??, January 2000. CODEN CCUJEX. ISSN 1075-2838.

Hanov:2000:LWW

[Han00] Steve Hanov. A lightweight window wrapper. C/C++ Users Journal, 18(8):26–??, August 2000. CODEN CCUJEX. ISSN 1075-2838.

Harmon:2002:PP

[Har02] Trevor Harmon. Probing the palm OS. C/C++ Users Journal, 20(2): ??, February 2002. CODEN CCU-JEX. ISSN 1075-2838.

Hicks:2000:CCT

[Hic00] Craig Hicks. C/C++ tips: Tip #1: Creating an index table in STL. C/C++ Users Journal, 18 (8):61–??, August 2000. CODEN CCUJEX. ISSN 1075-2838.

Howard:2000:UPW

[How00] David M. Howard. Using predicate waits with Win32 threads. C/C++ Users Journal, 18(5):18– ??, May 2000. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2002:MSS

[HR02] Moishe Halibard and Moshe Rubin. A multiple substring search algorithm. C/C++ Users Journal, 20(6):6–??, June 2002. CO-DEN CCUJEX. ISSN 1075-2838.

Hsiao:2000:SPF

[Hsi00] Gary Hsiao. A simple persistence framework. C/C++ Users Journal, 18(8):46–??, August 2000. CODEN CCUJEX. ISSN 1075-

Hubbard:2001:BPS

[Hub01a] John Hubbard. Building a professional software toolkit. C/C++ Users Journal, 19(5):42–??, May 2001. CODEN CCUJEX. ISSN 1075-2838.

Huber:2001:EFC

[Hub01b] Andreas Huber. Elegant function call wrappers. C/C++ Users Journal, 19(5):8–??, May 2001. CODEN CCUJEX. ISSN 1075-2838.

Hubbard:2002:BCR

[Hub02] John F. Hubbard. Binary code reuse in a Linux environment. C/C++ Users Journal, 20(3):??, March 2002. CODEN CCUJEX. ISSN 1075-2838.

Jarvi:2001:TTM

[Jär01] Jaakko Järvi. Tuple types and multiple return values. C/C++ Users Journal, 19(8):24–??, August 2001. CODEN CCUJEX. ISSN 1075-2838.

Kamper:2001:BRC

[Käm01] Stephan Kämper. Book review: Core C++ — A Software Engineering Approach. C/C++ Users Journal, 19(3):55–??, March 2001. CODEN CCUJEX. ISSN 1075-2838.

Karlsson:2002:SPB

[Kar02] Bjorn Karlsson. Smart pointers in Boost. C/C++ Users Journal, 20(4):34–??, April 2002. CODEN CCUJEX. ISSN 1075-2838.

Kelly-Bootle:2000:PMDe

[KB00a] Stan Kelly-Bootle. Post-mortem debunker: Angelology. C/C++ Users Journal, 18(10):88, October 2000. CODEN CCUJEX. ISSN 1075-2838.

Kelly-Bootle:2000:PMDc

[KB00b] Stan Kelly-Bootle. Post-mortem debunker: From MAW to MAM, more? C/C++ Users Journal, 18 (6):??, June 2000. CODEN CCU-JEX. ISSN 1075-2838.

Kelly-Bootle:2000:PMDa

[KB00c] Stan Kelly-Bootle. Post-mortem debunker: Pure vs. applied. C/C++ Users Journal, 18(2):104, February 2000. CODEN CCU-JEX. ISSN 1075-2838.

Kelly-Bootle:2000:PMDb

[KB00d] Stan Kelly-Bootle. Post-mortem debunker: Shall I compare thee … why not? C/C++ Users Journal, 18(4):96, April 2000. CODEN CCUJEX. ISSN 1075-2838.

Kelly-Bootle:2000:PMDd

[KB00e] Stan Kelly-Bootle. Post-mortem debunker: Whichever comes first. C/C++ Users Journal, 18(8):96, August 2000. CODEN CCUJEX. ISSN 1075-2838.

Kelly-Bootle:2001:PMDd

[KB01a] Stan Kelly-Bootle. Post-mortem debunker: Coder block. C/C++ Users Journal, 19(8):80–??, August 2001. CODEN CCUJEX. ISSN 1075-2838.

Kelly-Bootle:2001:PMDb

[KB01b] Stan Kelly-Bootle. Post-mortem debunker: Reply ALL. C/C++ Users Journal, 19(3):96–??, March 2001. CODEN CCUJEX. ISSN 1075-2838.

Kelly-Bootle:2001:PMDa

[KB01c] Stan Kelly-Bootle. Post-mortem debunker: Time flies. C/C++ Users Journal, 19(1):96–??, January 2001. CODEN CCUJEX. ISSN 1075-2838.

Kelly-Bootle:2001:PMDe

[KB01d] Stan Kelly-Bootle. Post-mortem debunker: Trust me. C/C++ Users Journal, 19(12):??, December 2001. CODEN CCUJEX. ISSN 1075-2838.

Kelly-Bootle:2001:PMDc

[KB01e] Stan Kelly-Bootle. Post-mortem debunker: What next? C/C++ Users Journal, 19(6):80–??, June 2001. CODEN CCUJEX. ISSN 1075-2838.

Kelly-Bootle:2002:PMD

[KB02] Stan Kelly-Bootle. Post-mortem debunker: Spurious precision. C/C++ Users Journal, 20(4):72, April 2002. CODEN CCUJEX. ISSN 1075-2838.

Kempf:2002:BTL

[Kem02a] Bill Kempf. The Boost.Threads library. C/C++ Users Journal, 20(5):6–??, May 2002. CODEN CCUJEX. ISSN 1075-2838.

Kempke:2002:ADB

[Kem02b] Christopher Kempke. Adaptable dialog boxes for cross-platform programming. C/C++ Users Journal, 20(8):13–14, 16–18, 20– 21, August 2002. CODEN CCU-JEX. ISSN 1075-2838.

Koss:2002:TDD

[KL02] Robert Koss and Jeff Langr. Test driven development in C/C++. C/C++ Users Journal, 20(10):6, 8, 10, 12–14, 16–17, October 2002. CODEN CCUJEX. ISSN 1075-2838.

Kleber:2000:TSA

[Kle00] Jeff Kleber. Thread-safe access to collections. C/C++ Users Journal, 18(5):36–??, May 2000. CO-DEN CCUJEX. ISSN 1075-2838.

Koenig:2001:CMEb

[KM01a] Andrew Koenig and Barbara E.

Moo. C++ made easier: Programs that work by accident. C/C++ Users Journal, 19(2):??, February 2001. CODEN CCU-JEX. ISSN 1075-2838.

Koenig:2001:CMEd

[KM01b] Andrew Koenig and Barbara E.

Moo. C++ made easier: The rule of three. C/C++ Users Journal, 19(6):65–??, June 2001. CODEN CCUJEX. ISSN 1075-2838.

Koenig:2001:CMEf

[KM01c] Andrew Koenig and Barbara E.

Moo. C++ made easier: Understanding binary searches. C/C++ Users Journal, 19(10):50–??, October 2001. CODEN CCUJEX. ISSN 1075-2838.

Koenig:2001:CMEc

[KM01d] Andrew Koenig and Barbara E.

Moo. C++ made easier: w vectors grow. C/C++ Users Journal, 19 (4):??, April 2001. CODEN CCU-JEX. ISSN 1075-2838.

Koenig:2001:CMEe

[KM01e] Andrew Koenig and Barbara E. Moo. C++ made easier: Which container should I use? C/C++ Users Journal, 19(8):58–??, August 2001. CODEN CCUJEX. ISSN 1075-2838.

Koenig:2002:CMEc

[KM02a] Andrew Koenig and Barbara E. Moo. C++ made easier: Avoiding array overruns. C/C++ Users Journal, 20(6):44–??, June 2002. CODEN CCUJEX. ISSN 1075-2838.

Koenig:2002:CMEd

[KM02b] Andrew Koenig and Barbara E. Moo. C++ made easier: Handles and exception safety, part 1: A simple handle class. C/C++ Users Journal, 20(8):32, 34–39, August 2002. CODEN CCUJEX. ISSN 1075-2838.

Koenig:2002:CMEe

[KM02c] Andrew Koenig and Barbara E. Moo. C++ made easier: Handles and exception safety, part 2: Intrusive reference counting. C/C++ Users Journal, 20(10): 45–50, October 2002. CODEN CCUJEX. ISSN 1075-2838.

Koenig:2002:CMEa

[KM02d] Andrew Koenig and Barbara E. Moo. C++ made easier: Naming unknown types. C/C++ Users Journal, 20(2):??, February 2002. CODEN CCUJEX. ISSN 1075-2838.

Koenig:2002:CMEb

[KM02e] Andrew Koenig and Barbara E.

Moo. C++ made easier: Plain old data. C/C++ Users Journal, 20(4):50–??, April 2002. CODEN CCUJEX. ISSN 1075-2838.

Koenig:2001:CMEa

[Koe01] Andrew Koenig. C++ made easier: Using library algorithms. C/C++ Users Journal, 19(1):42– ??, January 2001. CODEN CCU-JEX. ISSN 1075-2838.

Kougiouris:2000:SYC

[Kou00] Panos Kougiouris. Stressing your COM objects. C/C++ Users Journal, 18(4):30–??, April 2000. CODEN CCUJEX. ISSN 1075-2838.

Kuepper:2005:UFS

[Kue05] Tilman Kuepper. Unicode and filtering stream buffers. C/C++ Users Journal, 23(3):28, 30–33, March 2005. CODEN CCUJEX. ISSN 1075-2838.

Labbe:2000:SCP

[Lab00] Roger Labbe. Solving combinatorial problems with STL and backtracking. C/C++ Users Journal, 18(2):56–??, February 2000. CO-DEN CCUJEX. ISSN 1075-2838.

Lafreniere:2000:SMD

[Laf00] David Lafreniere. State machine design in C++. C/C++ Users Journal, 18(5):58–??, May 2000. CODEN CCUJEX. ISSN 1075-

Leman:2002:EFT

[Lem02] Dmitri Leman. An efficient and flexible tracing technique. C/C++ Users Journal, 20(4):24–??, April 2002. CODEN CCUJEX. ISSN 1075-2838.

Lingsong:2001:EDB

[Lin01] He Lingsong. Exchanging data between Java and C/C++ Windows programs. C/C++ Users Journal, 19(1):34–??, January 2001. CO-DEN CCUJEX. ISSN 1075-2838.

Lippman:2001:CD

[Lip01] Stanley B. Lippman. The C# delegate. C/C++ Users Journal, 19(7):42–??, July 2001. CODEN CCUJEX. ISSN 1075-2838.

Lippman:2002:NME

[Lip02] Stanley Lippman. The .NET managed extensions to C++. C/C++ Users Journal, 20(9):6, 8–11, September 2002. CODEN CCU-JEX. ISSN 1075-2838.

Liu:2001:IRM

[Liu01] Qiang Liu. Implementing reusable mathematical procedures using C++. C/C++ Users Journal, 19 (6):22–??, June 2001. CODEN CCUJEX. ISSN 1075-2838.

Lacoude:2000:PEC

[LK00] Philippe Lacoude and Grum Ketema. Programming Excel COM objects in C++. C/C++ Users Journal, 18(4):22–??, April 2000. CODEN CCUJEX. ISSN 1075-2838.

Lopiano:2001:SOA

[Lop01] Fabio Lopiano. Sending objects across platforms. C/C++ Users Journal, 19(10):10–??, October 2001. CODEN CCUJEX. ISSN 1075-2838.

Lorenzen:2001:AEI

[Lor01] Bob Lorenzen. An algorithm for efficient image warping. C/C++ Users Journal, 19(9):??, September 2001. CODEN CCUJEX. ISSN 1075-2838.

Loudon:2000:II

[Lou00] Kyle Loudon. An introduction to interpolation. C/C++ Users Journal, 18(2):20–??, February 2000. CODEN CCUJEX. ISSN 1075-2838.

Lowy:2000:MPO

[Low00] Juval Lowy. Making primitive objects thread safe. C/C++ Users Journal, 18(3):85–??, March 2000. CODEN CCUJEX. ISSN 1075-2838.

May:2001:ECC

[May01] Patrick May. Encapsulating CORBA components with the adapter and bridge patterns. C/C++ Users Journal, 19(2):??, February 2001. CODEN CCU-JEX. ISSN 1075-2838.

Meyers:2000:NCIc

[Mey00a] Randy Meyers. The new C: Integers in C99, part 1. C/C++ Users Journal, 18(12):65–??, December 2000. CODEN CCUJEX. ISSN 1075-2838.

Meyers:2000:NCIa

[Mey00b] Randy Meyers. The new C: Introducing C99. C/C++ Users Journal, 18(10):49–??, October 2000. CODEN CCUJEX. ISSN 1075-2838.

Meyers:2000:NCIb

[Mey00c] Randy Meyers. The new C: It all began with FORTRAN. C/C++ Users Journal, 18(11): 80–??, November 2000. CODEN CCUJEX. ISSN 1075-2838.

Meyers:2000:HNM

[Mey00d] Scott Meyers. How non-member functions improve encapsulation. C/C++ Users Journal, 18(2):44– ??, February 2000. CODEN CCU-JEX. ISSN 1075-2838.

Meyers:2001:NCC

[Mey01a] Randy Meyers. The new C: Compound literals. C/C++ Users Journal, 19(6):33–??, June 2001. CODEN CCUJEX. ISSN 1075-2838.

Meyers:2001:NCD

[Mey01b] Randy Meyers. The new C: Declarations and initializations. C/C++ Users Journal, 19(4):??, April 2001. CODEN CCUJEX. ISSN 1075-2838.

Meyers:2001:NCIa

[Mey01c] Randy Meyers. The new C: Integers, part 2. C/C++ Users Journal, 19(1):52–??, January 2001. CODEN CCUJEX. ISSN 1075-2838.

Meyers:2001:NCIb

[Mey01d] Randy Meyers. The new C: Integers, part 3. C/C++ Users Journal, 19(2):??, February 2001. CO-DEN CCUJEX. ISSN 1075-2838.

Meyers:2001:NCV

[Mey01e] Randy Meyers. The new C: Variable length arrays, part 2. C/C++ Users Journal, 19(12):??, December 2001. CODEN CCUJEX. ISSN 1075-2838.

Meyers:2001:NCW

[Mey01f] Randy Meyers. The new C: Why variable length arrays? C/C++ Users Journal, 19(10):46–??, October 2001. CODEN CCUJEX. ISSN 1075-2838.

Meyers:2001:NCX

[Mey01g] Randy Meyers. The new C: X macros. C/C++ Users Journal, 19(5):52–??, May 2001. CODEN CCUJEX. ISSN 1075-2838.

Meyers:2001:DSS

[Mey01h] Scott Meyers. Distinguishing STL search algorithms. C/C++ Users Journal, 19(12):??, December 2001. CODEN CCUJEX. ISSN 1075-2838.

Meyers:2001:SAV

[Mey01i] Scott Meyers. STL algorithms vs. hand-written loops. C/C++ Users Journal, 19(10):26–??, October 2001. CODEN CCUJEX. ISSN 1075-2838.

Meyers:2001:TGE

[Mey01j] Scott Meyers. Three guidelines for effective iterator usage. C/C++ Users Journal, 19(6):10–??, June 2001. CODEN CCUJEX. ISSN 1075-2838.

Meyers:2002:NCB

[Mey02a] Randy Meyers. The new C: bool, advice to C and C++ programmers. C/C++ Users Journal, 20 (5):–??, May 2002. CODEN CCU-JEX. ISSN 1075-2838.

Meyers:2002:NCI

[Mey02b] Randy Meyers. The new C: Inline functions. C/C++ Users Journal, 20(7):47–51, July 2002. CODEN CCUJEX. ISSN 1075-2838.

Meyers:2002:NCV

[Mey02c] Randy Meyers. The new C: VLAs, Part 4: VLA typedefs and flexible array members. C/C++ Users Journal, 20(3):??, March 2002. CODEN CCUJEX. ISSN 1075-2838.

Meyers:2002:VLA

[Mey02d] Randy Meyers. Variable length arrays, part 3: Pointers and parameters. C/C++ Users Journal, 20(1):??, January 2002. CODEN CCUJEX. ISSN 1075-2838.

Morana:2000:BBD

[Mor00a] Marco Morana. Browser-based directory access through LDAP and COM. C/C++ Users Journal, 18(7):34–??, July 2000. CODEN CCUJEX. ISSN 1075-2838.

Moreno:2000:HDT

[Mor00b] Carlos Moreno. HTML document templates for CGI applications. C/C++ Users Journal, 18 (9):10–??, September 2000. CO-DEN CCUJEX. ISSN 1075-2838.

Moreno:2001:WIC

[Mor01] Carlos Moreno. wxWindows, an intuitive cross-platform GUI framework. C/C++ Users Journal, 19(5):18–??, May 2001. CO-DEN CCUJEX. ISSN 1075-2838.

Morley:2002:UCT

[Mor02] Kevin T. Morley. Using constructed types in C++ unions. C/C++ Users Journal, 20(8):29– 31, August 2002. CODEN CCU-JEX. ISSN 1075-2838.

Mott:2002:WKA

[Mot02] Jason Mott. Writing KDE applications. C/C++ Users Journal, 20(3):??, March 2002. CODEN CCUJEX. ISSN 1075-2838.

Nakov:2001:AEM

[Nak01] Kalin Nakov. An asynchronous execution macro. C/C++ Users Journal, 19(11):??, November 2001. CODEN CCUJEX. ISSN 1075-2838.

Nash:2000:IFT

[Nas00] Alan Nash. Introduction to function try blocks. C/C++ Users Journal, 18(10):32–??, October 2000. CODEN CCUJEX. ISSN 1075-2838.

Nelson:2000:NPL

[Nel00] Erik Nelson. Network programming with Linux. C/C++ Users Journal, 18(9):26–??, September 2000. CODEN CCUJEX. ISSN 1075-2838.

Nelson:2001:CDS

[Nel01]  Mark Nelson.  A C++ de-

bug stream for Win32. C/C++ Users Journal, 19(9):??, September 2001. CODEN CCUJEX. ISSN 1075-2838.

Niec:2000:OSO

[Nie00] Todd Niec. Optimizing substring operations in string classes. C/C++ Users Journal, 18(10): 40–??, October 2000. CODEN CCUJEX. ISSN 1075-2838.

Nies:2002:TEE

[Nie02] Richard Nies. Tracing exceptions with an exception stack. C/C++ Users Journal, 20(4):14–??, April 2002. CODEN CCUJEX. ISSN 1075-2838.

Niklas:2000:CSD

[Nik00] Michal Niklas. A class for scanning directories. C/C++ Users Journal, 18(4):60–??, April 2000. CODEN CCUJEX. ISSN 1075-2838.

Nowak:2001:CCT

[Now01] Harald Nowak. C/C++ tips: Tip

#7: A remove if for vector<T

*>. C/C++ Users Journal, 19(7): 52–??, July 2001. CODEN CCU-JEX. ISSN 1075-2838.

Offer:2000:ELI

[Off00] Brad Offer. Error logging with Iostreams. C/C++ Users Journal, 18(4):52–??, April 2000. CODEN CCUJEX. ISSN 1075-2838.

Padamadan:2001:CMC

[Pad01] Babu George Padamadan. A cookie manager class for Web applications. C/C++ Users Journal, 19(8):10–??, August 2001. CO-DEN CCUJEX. ISSN 1075-2838.

Panzer:2001:TSC

[Pan01] John Panzer. A toolbox of string classes. C/C++ Users Journal, 19(7):10–??, July 2001. CODEN CCUJEX. ISSN 1075-2838.

Pee:2001:GWS

[Pee01] James Pee. Guidelines for wrapping sockets in classes. C/C++ Users Journal, 19(12):??, December 2001. CODEN CCUJEX. ISSN 1075-2838.

Pehrson:2001:DGI

[Peh01] Bret S. Pehrson. Documenting the graphic import filter interface. C/C++ Users Journal, 19(3):26– ??, March 2001. CODEN CCU-JEX. ISSN 1075-2838.

Perry:2000:RNSa

[Per00a] Michael L. Perry. A reusable nonlinear system solver, part 1. C/C++ Users Journal, 18(6):??, June 2000. CODEN CCUJEX. ISSN 1075-2838.

Perry:2000:RNSb

[Per00b] Michael L. Perry. A reusable nonlinear system solver, part 2. C/C++ Users Journal, 18(8):32– ??, August 2000. CODEN CCU-JEX. ISSN 1075-2838.

Peterson:2000:CCT

[Pet00] Mark Peterson. C/C++ tips:

Tip #4: Self destructing threads. C/C++ Users Journal, 18(12): 44–??, December 2000. CODEN CCUJEX. ISSN 1075-2838.

Philips:2000:DMS

[Phi00a] Lawrence Philips. The double metaphone search algorithm.

C/C++ Users Journal, 18(6):??, June 2000. CODEN CCUJEX. ISSN 1075-2838.

Phillips:2000:BRP

[Phi00b] Dwayne Phillips. Book review: The Pragmatic Programmer. C/C++ Users Journal, 18 (9):66–??, September 2000. CO-DEN CCUJEX. ISSN 1075-2838.

Phillips:2000:RIN

[Phi00c] Dwayne Phillips. Removing image noise with adaptive filters. C/C++ Users Journal, 18(12): 10–??, December 2000. CODEN CCUJEX. ISSN 1075-2838.

Phillips:2000:VDI

[Phi00d] Dwayne Phillips. Visualizing depth images via rendering. C/C++ Users Journal, 18(5):40– ??, May 2000. CODEN CCUJEX. ISSN 1075-2838.

Pinto:2002:PGC

[Pin02] Paulo Pinto. Programming in GNOME with C++. C/C++ Users Journal, 20(5):28–??, May 2002. CODEN CCUJEX. ISSN 1075-2838.

Plauger:2000:JAV

[Pla00a] P. J. Plauger. Just another version of Algol. C/C++ Users Journal, 18(1):??, January 2000. CODEN CCUJEX. ISSN 1075-2838.

Plauger:2000:SCCb

[Pla00b] P. J. Plauger. Standard C/C++: A singly linked list. C/C++ Users Journal, 18(2):8–??, February 2000. CODEN CCUJEX. ISSN 1075-2838.

Plauger:2000:SCCe

[Pla00c] P. J. Plauger. Standard C/C++:

exit(0). C/C++ Users Journal, 18(5):10, 14, 16, May 2000. CO-DEN CCUJEX. ISSN 1075-2838.

Plauger:2000:SCCa

[Pla00d] P. J. Plauger. Standard C/C++:

Java standard time. C/C++ Users Journal, 18(1):??, January 2000. CODEN CCUJEX. ISSN 1075-2838.

Plauger:2000:SCCd

[Pla00e] P. J. Plauger. Standard C/C++:

Testing C++ library conformance. C/C++ Users Journal, 18(4):10–??, April 2000. CODEN CCUJEX. ISSN 1075-2838.

Plauger:2000:SCCc

[Pla00f] P. J. Plauger. Standard C/C++:testing conformance. C/C++ Users Journal, 18(3):10–??, March 2000. CODEN CCUJEX. ISSN 1075-2838.

Plauger:2002:CCP

[Pla02] P. J. Plauger. The C/C++ programming language. C/C++ Users Journal, 20(10):22, 24, 26– 28, October 2002. CODEN CCU-JEX. ISSN 1075-2838.

Powell:2000:VNF

[PW00] Gary Powell and Martin Weiser.

Views, A new form of container adaptors. C/C++ Users Journal, 18(4):40–??, April 2000. CODEN CCUJEX. ISSN 1075-2838.

Pyritz:2002:TEL

[Pyr02] Bill Pyritz. A technique for efficient logging and monitoring of real-time systems. C/C++ Users Journal, 20(3):??, March 2002. CODEN CCUJEX. ISSN 1075-2838.

Queisser:2000:CRW

[Que00] Andrew Queisser. Classes for reading and writing parameter blocks. C/C++ Users Journal, 18(10):22–??, October 2000. CO-DEN CCUJEX. ISSN 1075-2838.

Qureshi:2001:IRU

[Qur01] Shehrzad Qureshi. Image rotation using OpenGL texture maps. C/C++ Users Journal, 19(9):??, September 2001. CODEN CCU-JEX. ISSN 1075-2838.

Reichard:2000:SC

[Rei00] Douglas Reichard. Shim classes. C/C++ Users Journal, 18(2):53– ??, February 2000. CODEN CCU-JEX. ISSN 1075-2838.

Ring:2001:MPA

[Rin01] Michael C. Ring. MAPM, A portable arbitrary precision math library in C. C/C++ Users Journal, 19(11):??, November 2001. CODEN CCUJEX. ISSN 1075-2838.

Rogers:2000:CFC

[Rog00a] Joey Rogers. Copying files from the clipboard to a command prompt. C/C++ Users Journal, 18(7):28–??, July 2000. CODEN CCUJEX. ISSN 1075-2838.

Rogers:2000:SON

[Rog00b] Joey Rogers. The self-organizing neural network. C/C++ Users Journal, 18(6):??, June 2000. CO-DEN CCUJEX. ISSN 1075-2838.

Rogers:2002:EVM

[Rog02] Joey Rogers. Encapsulating virtual memory in Windows. C/C++ Users Journal, 20(1):??, January 2002. CODEN CCUJEX. ISSN 1075-2838.

Saks:2000:CTPa

[Sak00a] Dan Saks. C++ theory and practice: Replacing character arrays with strings, part 1. C/C++ Users Journal, 18(1):??, January 2000. CODEN CCUJEX. ISSN 1075-2838.

Saks:2000:CTPb

[Sak00b] Dan Saks. C++ theory and practice: Replacing character arrays with strings, part 2. C/C++ Users Journal, 18(2):65– ??, February 2000. CODEN CCU-JEX. ISSN 1075-2838.

Scanlon:2000:SWB

[Sca00] Mike Scanlon. Secure Web-based licensing. C/C++ Users Journal, 18(11):42–??, November 2000. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:2000:UEA

[Sch00a] Bobby Schmidt. Uncaught exceptions: Adventures of Aquaman. C/C++ Users Journal, 18(7):85– ??, July 2000. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:2000:UEC

[Sch00b] Bobby Schmidt. Uncaught excep-

tions: comp.cuj.qa.moderated.

C/C++ Users Journal, 18(3):89– ??, March 2000. CODEN CCU-JEX. ISSN 1075-2838.

Schmidt:2000:UED

[Sch00c] Bobby Schmidt. Uncaught exceptions: Dies irae. C/C++ Users Journal, 18(2):85–??, February 2000. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:2000:UEN

[Sch00d] Bobby Schmidt. Uncaught exceptions: No no .NET. C/C++ Users Journal, 18(12):73–??, December 2000. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:2000:UER

[Sch00e] Bobby Schmidt. Uncaught exceptions: Remembrance. C/C++ Users Journal, 18(5):75–??, May 2000. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:2000:UESb

[Sch00f] Bobby Schmidt. Uncaught exceptions: Semper fi. C/C++ Users Journal, 18(6):??, June 2000. CO-DEN CCUJEX. ISSN 1075-2838.

Schmidt:2000:UESc

[Sch00g] Bobby Schmidt. Uncaught exceptions: Soft geek cafe. C/C++ Users Journal, 18(8):83–??, August 2000. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:2000:UESa

[Sch00h] Bobby Schmidt. Uncaught exceptions: Standard deviation. C/C++ Users Journal, 18(4):81– ??, April 2000. CODEN CCU-JEX. ISSN 1075-2838.

Schmidt:2000:UESd

[Sch00i] Bobby Schmidt. Uncaught exceptions: Sunshine superman.

C/C++ Users Journal, 18(10): 73–??, October 2000. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:2000:UET

[Sch00j] Bobby Schmidt. Uncaught exceptions: Tempus fugit. C/C++ Users Journal, 18(1):??, January 2000. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:2000:UEB

[Sch00k] Bobby Schmidt. Uncaught exceptions: The ballad of Ned Kelly. C/C++ Users Journal, 18(9):94– ??, September 2000. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:2001:UEC

[Sch01a] Bobby Schmidt. Uncaught exceptions: 21st Century Man. C/C++ Users Journal, 19(1):80–??, January 2001. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:2001:UEAb

[Sch01b] Bobby Schmidt. Uncaught exceptions: All good things. C/C++ Users Journal, 19(6):56–??, June 2001. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:2001:UEAa

[Sch01c] Bobby Schmidt. Uncaught exceptions: An Alan Smithee film. C/C++ Users Journal, 19(3):80– ??, March 2001. CODEN CCU-JEX. ISSN 1075-2838.

Schmidt:2001:UEE

[Sch01d] Bobby Schmidt. Uncaught exceptions: eWriter. C/C++ Users Journal, 19(4):??, April 2001. CO-DEN CCUJEX. ISSN 1075-2838.

Schmidt:2001:UEF

[Sch01e] Bobby Schmidt. Uncaught exceptions: Fuzzy math. C/C++ Users Journal, 19(2):??, February 2001. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:2001:UEI

[Sch01f] Bobby Schmidt. Uncaught exceptions: International experts of mystery. C/C++ Users Journal, 19(9):??, September 2001. CO-DEN CCUJEX. ISSN 1075-2838.

Schmidt:2001:UEN

[Sch01g] Bobby Schmidt. Uncaught exceptions: Nevermind. C/C++ Users Journal, 19(8):65–??, August 2001. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:2001:UES

[Sch01h] Bobby Schmidt. Uncaught exceptions: Standard languages: Ask for them by name! C/C++ Users Journal, 19(10):38–??, October 2001. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:2001:UEG

[Sch01i] Bobby Schmidt. Uncaught exceptions: The great panjandrum. C/C++ Users Journal, 19(5):73– ??, May 2001. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:2001:UEY

[Sch01j] Bobby Schmidt. Uncaught exceptions: Your witness. C/C++ Users Journal, 19(11):??, November 2001. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:2002:STN

[Sch02a] Bobby Schmidt. Straight talk on the name and content of Redmond’s compilers and why

operator bool() is rarely the

right choice. C/C++ Users Journal, 20(9):42, 44–48, September 2002. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:2002:UED

[Sch02b] Bobby Schmidt. Uncaught exceptions: Don’t read this title! C/C++ Users Journal, 20(5):50– ??, May 2002. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:2002:UER

[Sch02c] Bobby Schmidt. Uncaught exceptions: Retreat. C/C++ Users Journal, 20(1):??, January 2002. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:2002:UESa

[Sch02d] Bobby Schmidt. Uncaught exceptions: Space-time discontinuum. C/C++ Users Journal, 20(3):??, March 2002. CODEN CCUJEX. ISSN 1075-2838.

Schmidt:2002:UESb

[Sch02e] Bobby Schmidt. Uncaught exceptions: The spiral dance. C/C++ Users Journal, 20(7):52, 54–56, 58, July 2002. CODEN CCUJEX. ISSN 1075-2838.

Seward:2002:SCC

[Sew02] Robert Y. Seward. A stream class for calling perl from C++. C/C++ Users Journal, 20(1):??, January 2002. CODEN CCUJEX. ISSN 1075-2838.

Smereka:2000:TIS

[Sme00] Richard Smereka. A TCP/IP socket location server. C/C++ Users Journal, 18(9):34–??, September 2000. CODEN CCUJEX. ISSN 1075-2838.

Smereka:2001:TCS

[Sme01] Richard Smereka. Two classes for simple socket programming. C/C++ Users Journal, 19(5):34– ??, May 2001. CODEN CCUJEX. ISSN 1075-2838.

Smith:2000:SBW

[Smi00] Mark L. Smith. An STL-based Nway set. C/C++ Users Journal, 18(3):76–??, March 2000. CODEN CCUJEX. ISSN 1075-2838.

Smith:2002:IND

[Smi02] J. Daniel Smith. Inside .NET’s delegates and events using C++. C/C++ Users Journal, 20(9):12, 14, 16–18, 20–21, September 2002. CODEN CCUJEX. ISSN 1075-2838.

Sorfa:2001:OSI

[Sor01] Petr Sorfa. Open source IDEs for Linux/Unix. C/C++ Users Journal, 19(3):8–??, March 2001. CO-DEN CCUJEX. ISSN 1075-2838.

Stanley:2001:AET

[Sta01] Ben Stanley. Adding exception testing to unit tests. C/C++ Users Journal, 19(4):??, April 2001. CODEN CCUJEX. ISSN 1075-2838.

Stroustrup:2002:CCCb

[Str02a] Bjarne Stroustrup. C and C++: A case for compatibility. C/C++ Users Journal, 20(8):22, 24–28, August 2002. CODEN CCUJEX. ISSN 1075-2838.

Stroustrup:2002:CCC

[Str02b] Bjarne Stroustrup. C and C++:

Case studies in compatibility. C/C++ Users Journal, 20(9):22, 24–28, 30–31, September 2002. CODEN CCUJEX. ISSN 1075-2838.

Stroustrup:2002:CCS

[Str02c] Bjarne Stroustrup. C and C++:

Siblings. C/C++ Users Journal, 20(7):28, 30, 32–36, July 2002. CODEN CCUJEX. ISSN 1075-2838.

Sutter:2000:SMC

[Sut00] Herb Sutter. Sutter’s Mill: Constructor failures (or, the objects that never were). C/C++ Users Journal, 18(11):72–??, November 2000. CODEN CCUJEX. ISSN 1075-2838.

Sutter:2001:CCR

[Sut01a] Herb Sutter. C++ conformance roundup. C/C++ Users Journal, 19(4):??, April 2001. CODEN CCUJEX. ISSN 1075-2838.

Sutter:2001:SMC

[Sut01b] Herb Sutter. Sutter’s Mill: Containers in memory: How big is big? C/C++ Users Journal, 19 (1):58–??, January 2001. CODEN CCUJEX. ISSN 1075-2838.

Sutter:2001:SMS

[Sut01c] Herb Sutter. Sutter’s Mill: The string formatters of Manor Farm. C/C++ Users Journal, 19(11):??, November 2001. CODEN CCU-JEX. ISSN 1075-2838.

Sutter:2001:SMN

[Sut01d] Herb Sutter. Sutter’s Mill: To new, perchance to throw (part 2 of 2). C/C++ Users Journal, 19(5): 66–??, May 2001. CODEN CCU-JEX. ISSN 1075-2838.

Sutter:2001:SMV

[Sut01e] Herb Sutter. Sutter’s Mill: Virtuality. C/C++ Users Journal, 19 (9):??, September 2001. CODEN CCUJEX. ISSN 1075-2838.

Sutter:2001:SMW

[Sut01f] Herb Sutter. Sutter’s Mill: Why not specialize function templates? C/C++ Users Journal, 19(7):65– ??, July 2001. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:2002:MCD

[Sut02a] Herb Sutter. Book review: Modern C++ Design. C/C++ Users Journal, 20(4):41–??, April 2002. CODEN CCUJEX. ISSN 1075-2838.

Sutter:2002:ETI

[Sut02b] Herb Sutter. Extensible templates: Via inheritance or traits? C/C++ Users Journal, 20(2):??, February 2002. CODEN CCU-JEX. ISSN 1075-2838.

Sutter:2002:SME

[Sut02c] Herb Sutter. Sutter’s Mill — “export” restrictions, part 1. C/C++ Users Journal, 20(9):50, 52–55, September 2002. CODEN CCU-JEX. ISSN 1075-2838.

Sutter:2002:SMP

[Sut02d] Herb Sutter. Sutter’s Mill: A pragmatic look at exception specifications. C/C++ Users Journal, 20(7):59–62, 64, July 2002. CO-DEN CCUJEX. ISSN 1075-2838.

Sutter:2002:MHS

[Sut02e] Herb Sutter. Sutter’s Mill: The migrating habits of the Standard C++ library: a work in progress. C/C++ Users Journal, 20(5):42– ??, May 2002. CODEN CCUJEX. ISSN 1075-2838.

Sutter:2002:SMT

[Sut02f] Herb Sutter. Sutter’s Mill: Toward a Standard C++0x library, part 1. C/C++ Users Journal, 20(1):??, January 2002. CODEN CCUJEX. ISSN 1075-2838.

Swaim:2000:SLR

[Swa00] David C. Swaim II. A simple linear regression class. C/C++ Users Journal, 18(8):62–??, August 2000. CODEN CCUJEX. ISSN 1075-2838.

Telford:2000:TDP

[Tel00] Alex Telford. Taming the 3-D perspective transform. C/C++ Users Journal, 18(3):57–??, March 2000. CODEN CCUJEX. ISSN 1075-2838.

Tennberg:2002:RGO

[Ten02] Patrick Tennberg. Refactoring global objects in multithreaded applications. C/C++ Users Journal, 20(5):20–??, May 2002. CO-DEN CCUJEX. ISSN 1075-2838.

Thierauf:2000:EDP

[Thi00] Rainer Thierauf. Extracting data from X-Y plots. C/C++ Users Journal, 18(11):10–??, November 2000. CODEN CCUJEX. ISSN 1075-2838.

Tilevich:2001:EST

[Til01] Eli Tilevich. Extending the standard template library with association classes. C/C++ Users Journal, 19(9):??, September 2001. CODEN CCUJEX. ISSN 1075-2838.

Timperley:2001:CCT

[Tim01] Mark Timperley. C/C++ tips: Tip #5: A C/C++ comment macro. C/C++ Users Journal, 19 (1):41–??, January 2001. CODEN CCUJEX. ISSN 1075-2838.

Trudell:2002:AWC

[Tru02] Bill Trudell. The Application Watchman class. C/C++ Users Journal, 20(7):16, 18–20, 22–24, 26–27, July 2002. CODEN CCU-JEX. ISSN 1075-2838.

Vlasceanu:2001:GCB

[Vla01] Cristian Vlasceanu. Generalizing the concepts behind auto ptr. C/C++ Users Journal, 19(8):36– ??, August 2001. CODEN CCU-JEX. ISSN 1075-2838.

Vlasceanu:2002:DME

[Vla02] Cristian Vlasceanu. Debugging memory errors with custom allocators. C/C++ Users Journal, 20(4):6–??, April 2002. CODEN CCUJEX. ISSN 1075-2838.

Wall:2001:BRR

[Wal01] Graham Wall. Book review: A Real-Time Programmer’s Review of mC/OS-II. C/C++ Users Journal, 19(10):32–??, October 2001. CODEN CCUJEX. ISSN 1075-2838.

Weihe:2002:TIS

[Wei02] Karsten Weihe. Towards improved static safety: Expressing meaning by type. C/C++ Users Journal, 20(9):32, 34–36, 38–40, September 2002. CODEN CCUJEX. ISSN 1075-2838.

Wu:2000:ECC

[Wu00] Wenfei Wu. Effective contour creation with OpenGL texture mapping. C/C++ Users Journal, 18 (12):24–??, December 2000. CO-DEN CCUJEX. ISSN 1075-2838.

Yoder:2001:UTF

[Yod01] Michael B. Yoder. Using template functions to customize library behavior. C/C++ Users Journal, 19 (3):42–??, March 2001. CODEN CCUJEX. ISSN 1075-2838.

York:2001:RAC

[Yor01] Kyle York. A random access compressed file layer. C/C++ Users Journal, 19(7):32–??, July 2001. CODEN CCUJEX. ISSN 1075-2838.

Zack:2001:DUG

[Zac01] Randy Zack. Debugging under GNU/Linux. C/C++ Users Journal, 19(2):??, February 2001. CO-DEN CCUJEX. ISSN 1075-2838.

Zolman:2001:SEM

[Zol01] Leor Zolman. An STL error message decryptor for Visual C++. C/C++ Users Journal, 19(7):24– ??, July 2001. CODEN CCUJEX. ISSN 1075-2838.